### v1

# A Mechanistic Interpretation of Syllogistic Reasoning in Auto-Regressive Language Models

Abstract
Recent studies on logical reasoning in auto-regressive Language Models (LMs) have sparked a debate on whether such models can learn systematic reasoning principles during pre-training or merely exploit superficial patterns in the training data. This paper presents a mechanistic interpretation of syllogistic reasoning in LMs to further enhance our understanding of internal dynamics. Specifically, we present a methodology for circuit discovery aimed at disentangling content-independent reasoning mechanisms from world knowledge acquired during pre-training. Through two distinct intervention methods, we uncover a sufficient and necessary circuit involving middle-term suppression that elucidates how LMs transfer information to derive valid conclusions from premises. Furthermore, we investigate how belief biases manifest in syllogistic reasoning, finding evidence of partial contamination from additional attention heads responsible for encoding commonsense and contextualized knowledge. Finally, we explore the generalization of the discovered mechanisms across various syllogistic schemes and model sizes, finding that the identified circuit is sufficient and necessary for all the schemes on which the model achieves high downstream accuracy ($\geq$ 60%). Overall, our findings suggest that LMs indeed learn transferable content-independent reasoning mechanisms, but that, at the same time, such mechanisms do not involve generalisable and abstract logical primitives, being susceptible to contamination by the same world knowledge acquired during pre-training.
Introduction
Transformer-based Language Models (LMs) have led to remarkable results across various natural language processing tasks (Radford et al. 2018, 2019; Brown et al. 2020; OpenAI et al. 2024; Jason et al. 2022; Bubeck et al. 2023). This success has catalyzed research interest in systematically exploring the reasoning capabilities emerging during pre-training (Clark, Tafjord, and Richardson 2020). Recent findings suggest that logical reasoning abilities may emerge in large-scale models (Rae et al. 2022; Kojima et al. 2022; Wei et al. 2022) or through transfer learning on specialized datasets (Betz, Voigt, and Richardson 2021). However, ongoing debates question whether these models apply systematic inference rules or reuse superficial patterns learned during pre-training (Talmor et al. 2020; Kassner, Krojer, and Schütze 2020; Wu et al. 2024). This controversy underscores the need for a deeper understanding of the low-level logical inference mechanisms in LMs (Rozanova, Valentino, and Freitas 2024; Rozanova et al. 2023, 2022).
Figure 1: Conceptual representation of the circuit for processing symbolic syllogisms: (Long Induction) Early layers exhibit biases towards the wrong conclusion due to long-range repetition of the first premise “All A are B”. (Duplication) Induction heads aggregate information about duplicated middle terms. (Suppression) The model inhibits middle-term information (i.e., ‘B’), suppressing the long induction mechanism. (Mover) Token-specific information is propagated to the last token position. The process culminates in the prediction shift from ‘B’ to the correct token, ‘C.’
To improve our understanding of the internal mechanisms, this paper focuses on Mechanistic Interpretability (Olah et al. 2020; Nanda et al. 2023), aiming to discover the core circuit responsible for content-independent syllogistic reasoning. In particular, we focus on categorical syllogisms with universal affirmative quantifiers (i.e., AAA-1, Barbara) motivated by two key factors. First, as observed in natural logic studies (MacCartney and Manning 2007), this form of syllogistic reasoning is prevalent in everyday language. Therefore, it is likely that LMs are exposed to such reasoning schema during pre-training. Second, AAA-1 is a form of unconditionally valid syllogism independent of the premises’ truth condition (Holyoak and Morrison 2005). This characteristic offers a deterministic and scalable task design, other than allowing us to investigate the disentanglement between reasoning and knowledge representation.
Through the adoption of Activation Patching (Meng et al. 2022) and embedding space analysis (i.e., Logit Lens) (Nostalgebraist 2020; Geva et al. 2022; Dar et al. 2023), we investigate the following main research questions: RQ1: What is the content-independent syllogistic reasoning mechanism internalized in LMs during pre-training?; RQ2: Are content-independent mechanisms disentangled from specific world knowledge and belief biases?; RQ3: Does the core reasoning mechanism generalize across model sizes and different syllogistic schemes?
To answer these questions, we present a methodology that consists of 3 main stages. First, we define a syllogistic completion task designed to assess the model’s ability to predict valid conclusions from premises and facilitate the construction of test sets for circuit analysis. Second, we implement a circuit discovery pipeline on the syllogistic schema instantiated only with symbolic variables (Table 1, Symbolic) to identify the core sub-components responsible for content-independent reasoning. We conduct this analysis under two intervention methods: middle-term corruption and all-term corruption, aiming to identify latent transitive reasoning mechanisms and term-related information flow. Third, we investigate the generalization of the identified circuit on concrete schemes instantiated with commonsense knowledge to identify potential belief biases and explore how the internal behavior varies with different schemes and model sizes.
We present the following overall conclusions:
  1. 1.
    The circuit analysis reveals that auto-regressive LMs develop specific inference mechanisms during pre-training, finding evidence supporting a three-stage mechanism for syllogistic reasoning: (1) naive recitation of the first premise; (2) suppression of duplicated middle-term information; and (3) mediation towards the correct output through the interplay of attention heads.
  2. 2.
    Further experiments on circuit transferability demonstrate that the identified mechanism is still necessary for reasoning on syllogistic schemes instantiated with commonsense knowledge. However, a deeper analysis suggests that specific belief biases acquired during pre-training might play an important role, contaminating the content-independent circuit mechanism with additional attention heads responsible for encoding general and contextualized world knowledge.
  3. 3.
    We found that the identified circuit is sufficient and necessary for all the syllogistic schemes in which the model achieves high downstream accuracy ($\geq$ 60%). This result suggests that LMs learn content-independent reasoning mechanisms transferable across abstract schemes.
  4. 4.
    The intervention results on models with different sizes show similar suppression mechanism patterns and information flow. However, we found evidence that the contribution of attention heads becomes more complex with model sizes, further supporting the hypothesis of increasing contamination from external world knowledge.
Our following study is conducted using TransformerLens (Nanda and Bloom 2022) on an Nvidia A100 GPU with 80GB of memory. The dataset and code to reproduce our experiments are available online to encourage future work in the field [^1] http://anonynous˙url.com.
Methodology
Our main research objective is to discover and interpret the core mechanisms adopted by auto-regressive Language Models (LMs) when performing content-independent syllogistic reasoning. To this end, we present a methodology that consists of 3 main stages. First, we define a syllogistic completion task that can be instrumental for our analysis. Second, we leverage the syllogistic completion task to implement a circuit discovery pipeline on a syllogistic schema instantiated only with symbolic variables (Table 1, Symbolic). Third, we investigate the generalization of the identified circuit on concrete schemes instantiated with commonsense knowledge and explore how the internal behavior varies with different schemes and model sizes.
Premises ($\mathcal{P}_{1},\mathcal{P}_{2}$) Conclusion ($\mathcal{C}$)
Symbolic
All A are B.
All B are C.
Therefore,
all A are C.
Belief-consistent (True premises)
All men are humans.
All humans are mortal.
Therefore,
all men are mortal.
Belief-inconsistent (False premises)
All pilots are people.
All people are blond.
Therefore,
all pilots are blond.
Table 1: Examples of the same syllogistic schema (i.e., AAA-1, Barbara) instantiated with different content (with the expected completion in bold). The ability to complete the syllogism is a function of the reasoning schema, independent of the specific instantiation or truth condition of the premises.
Figure 2: The conceptual pipeline of the symbolic circuit analysis on the syllogism completion task. Figures (a)–(e) illustrate the stages involved in processing a syllogistic argument: argument formalization (a), construction of clean (b) and corrupted inputs (c), activation patching (d) and embedding space analysis (i.e., Logit Lens) (e). The uncovered symbolic circuit is subsequently evaluated to determine whether the core reasoning schema operates independently from world knowledge or belief bias.
Syllogism Completion Task
We design a syllogism completion task for assessing the reasoning abilities of LMs, building upon established approaches in the literature (Betz, Voigt, and Richardson 2021; Wu et al. 2023). In particular, we focus on categorical syllogisms with universal affirmative quantifiers (i.e., AAA-1, Barbara) because of their frequency in natural language and their content-independent reasoning property.
The syllogistic argument is typically composed of two premises ($\mathcal{P}_{1},\mathcal{P}_{2}$) and a conclusion ($\mathcal{C}$), with $s$ and $p$ denoting the subject and predicate terms in the conclusion (e.g. men and mortal), and $m_{1}$ and $m_{2}$ (e.g., humans) denoting the middle terms in the two premises, with $m_{1}\equiv m_{2}$ in the case of the AAA-1 syllogism.
To evaluate syllogistic reasoning in LMs, we formalize the completion task as masking language modeling, removing the final token $p$ from the conclusion (e.g., All men are humans. All humans are mortal. Therefore, all men are), and comparing the probability assigned by the LM to $p$ (e.g., mortal) and the middle term $m$ (e.g., humans). In general, an LM is successful in the completion of a task if the following condition applies:
$P(p\mid[\mathcal{P}_{1};\mathcal{P}_{2};\mathcal{C}\setminus\{p\}])>P(m\mid[\mathcal{P}_{1};\mathcal{P}_{2};\mathcal{C}\setminus\{p\}])$
In our experiments, we measure the logit difference $\delta$ between the tokens $p$ and $m$, defined as $\delta(p,m)=\text{logit}(p)-\text{logit}(m)$, which approximates the log ratio of the conditional probabilities:
$\delta(p,m)\approx\log{\left(\frac{P(p\mid[\mathcal{P}_{1};\mathcal{P}_{2};\mathcal{C}\setminus\{p\}])}{P(m\mid[\mathcal{P}_{1};\mathcal{P}_{2};\mathcal{C}\setminus\{p\}])}\right)}$
Dataset Construction.
We conduct experiments using two distinct datasets, symbolic and non-symbolic, to derive comparative implications for model reasoning capabilities. The symbolic dataset is constructed by randomly sampling triples (e.g., A, B, and C) from the set of 26 uppercase letters of the English alphabet. On the other side, the non-symbolic dataset is constructed by replacing the letters with words while preserving the syllogistic schema. Here, we generate two different non-symbolic sets: a belief-consistent set containing true premises and a belief-inconsistent set containing false premises (see Table 1). To guarantee the truth condition of the premises, we leverage GenericsKB (Bhakthavatsalam, Anastasiades, and Clark 2020), a knowledge base containing statements about commonsense knowledge. The detailed generation process is in Appendix B.
Circuit Discovery
Our main objective is to find a circuit $\mathcal{C}$ for syllogistic reasoning in a Language Model $\mathcal{M}$. A circuit $\mathcal{C}$ can be defined as a subset of the original model $\mathcal{M}$ that is both sufficient and necessary for achieving the original model performance on the syllogistic completion task. In order to identify a circuit, we employ Activation Patching together with circuit ablation methods (Meng et al. 2022; Vig et al. 2020).
Activation Patching.
This technique involves modifying the activation of targeted components and observing the resulting changes. Our study primarily examines the activation of residual streams and multi-head self-attention at the sequence level to trace the term information flow. Given a masked syllogistic input $x=[\mathcal{P}_{1};\mathcal{P}_{2};\mathcal{C}\setminus\{p\}]$ as $(s,m_{1},m_{2},p)$, which can be read as “All $s$ are $m_{1}$. All $m_{2}$ are $p$. Therefore, all $s$ are [mask]”, and its correct completion $y=p$, the Activation Patching technique consists of the following steps:
  1. (1)
    Clean Run: For the target input $(s,m_{1},m_{2},p)$, we record the baseline logit difference, $\delta_{+}(p,m)$ from the forward pass output of the model (Figure 2(1)).
  2. (2)
    Corrupted Run: We re-run the model on a corrupted input after applying a specific intervention (e.g., changing the middle term $m_{2}$) and record the logit difference $\delta_{-}(p,m)$ (Figure 2(3)).
  3. (3)
    Intervened Run: We run the model with the corrupted inputs again and replace the corrupted activations with those from the clean runs to compute the response from the remaining components and measure the causal impact of the intervention. Here, we measure the adjusted logit difference $\delta_{p}(p,m)$ (Figure 2(2)).
  4. (4)
    Quantification: We quantify the causal impact of each intervention using a patching score $\mathcal{S}$ following (Heimersheim and Nanda 2024) as shown in Figure 2(4). This score is further normalized to $[-1,1]$.
We complement Activation Patching with known methods for analysing hidden activations in transformers, including Logit Lens (Nostalgebraist 2020) with an input-agnostic approach (Dar et al. 2023; Hanna, Liu, and Variengien 2024). Additional technical details can be found in Appendix C.
Interventions.
The choice of tokens to corrupt is a critical aspect of the experimental design, and it is essential to establish an appropriate type of intervention that aligns with the hypothesis being tested. In this study, we employ two distinct interventions to isolate the mechanisms related to the reasoning schema from the propagation of specific token information:
  1. (1)
    Middle-Term Corruption: To investigate the reasoning mechanism employed in the syllogism completion task, we primarily focus on the transitive property of the middle term, $(m_{1}\rightarrow m_{2})\wedge(m_{1}\equiv m_{2})$. We hypothesize that disrupting the transitive property will localize the component responsible for syllogistic reasoning. Therefore, our intervention method replaces the second middle term $m_{2}$ with an unseen symbol $m_{2}^{\prime}$, breaking the equality and effectively corrupting the validity of the reasoning, pivoting the correct answer towards $m$: $(s,m_{1},m_{2},p)\rightarrow(s,m_{1},m_{2}^{\prime},p)$.
  2. (2)
    All-Term Corruption: We examine how term-related information flows to the last position for the final prediction to identify potential mover heads. To this end, we replace the original terms $(s,m_{1},m_{2},p)$ with different terms $(s^{\prime},m_{1}^{\prime},m_{2}^{\prime},p^{\prime})$ while keeping the completion answers unchanged. We hypothesize that if the mover heads carry the $p$ related information over $m$ related information for a valid prediction, the logit difference will increase; otherwise, it will decrease.
Figure 3: Accuracy on the syllogism completion task across all scales of GPT-2. We observe a phase change from small to medium models in overall performance.
Circuit Ablation.
Once we discover a relevant circuit $\mathcal{C}$, we evaluate its necessity and sufficiency using mean ablation. Specifically, we define a circuit $\mathcal{C}$ as necessary if ablating the identified heads $\mathcal{H}$ in $\mathcal{C}$ and preserving the remaining components in $\mathcal{M}$ decreases the original performance. Conversely, we define a circuit $\mathcal{C}$ as sufficient if preserving only the identified heads $\mathcal{H}$ in $\mathcal{C}$ and ablating all the remaining heads in the original model $\mathcal{M}$ is sufficient to obtain the performance of $\mathcal{M}$. To perform mean ablation, we gradually average the target attention heads from downstream to upstream layers and measure the average logit difference $\delta(p,m)$ to assess the impact of the ablation on $\mathcal{M}$. Additional details are included in Appendix D.
Empirical Evaluation
Figure 4: Comprehensive results of the symbolic circuit analysis. (a) Attention output patching results and (b) residual stream patching results in the middle-term intervention setup. (c) OV circuit logit lens results for $\mathbf{h}_{11.10}$, with input and output comprising 26 uppercase letters. (d) Residual stream patching results in the all-term corruption setup. (e)-(h) Average attention patching weights across the batch: (e) weights for token position $[p]$ attending, and (f)-(h) weights for the last token position attending. For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.
Model Selection.
We select the LM for the circuit analysis based on a trade-off between model size and performance. To this end, we measure the accuracy of GPT-2 (Radford et al. 2019) with different sizes (117M, 345M, 762M, and 1.5B) on the syllogism completion task. Specifically, we aim to identify potential phase transition points where significant performance changes emerge (Jason et al. 2022; Kaplan et al. 2020). As shown in Figure 3, we observe a notable phase transition from small to medium sizes, where average performance increases by 139.97% in the symbolic setting and 119.21% in the non-symbolic setting. Given these results, we select GPT-2 Medium for circuit discovery and subsequently test the compatibility with different model sizes (See Appendix E for the model architecture). For all experiments, we use 90 samples each for symbolic and non-symbolic arguments to balance the depth of analysis with computational efficiency.
Localization of Transitive Reasoning Mechanisms
Middle-term corruption reveals information flow relevant to the transitive property.
Figure 4(a-b) presents the results of the middle-term intervention targeting attention heads and residual stream. Figure 4(a), in particular, reveals the positive role of heads $\mathbf{h}_{11.10}$ and $\mathbf{h}_{19.1}$ (where $\mathbf{h}_{l.k}$ denotes the $k$th head in layer $l$). Moreover, we observe an information propagation pattern from the $[m_{2}]$ position to the $[p]$ position (Figure 4(b)). These observations allow us to hypothesize that information from $[m_{2}]$ is conveyed to $[p]$ on the residual stream by attention head $\mathbf{h}_{11.10}$, which exhibits a strong patching score around the layer responsible for the propagation. To verify this hypothesis, we further investigate the attention weights between $[p]$ and $[m_{2}]$. As expected, $[m_{2}]$ is the position most highly attended by $[p]$, with an average attention weight of $0.15\pm 0.07$ (Figure 4(e)). These results suggest that information is indeed moved from $[m_{2}]$ to $[p]$ on the residual stream subspace, with $\mathbf{h}_{11.10}$ playing a crucial role in this mechanism.
Duplicate middle-term information is aggregated via induction heads.
We further investigate how information from $[m_{2}]$ is moved to $[p]$, positing that this relates to the model’s internal reasoning mechanism. We notice that at position $[p]$, the model can observe the complete AAA-1 syllogistic structure (Figure 2(a)) with middle-term duplication. We hypothesize that this structural information for reasoning is collected at one position, given that information refinement occurs at the specific position such as last token (Hanna, Liu, and Variengien 2024; Stolfo, Belinkov, and Sachan 2023). To verify this, we trace the information flow of head $\mathbf{h}_{11.10}$ using path patching (Wang et al. 2023). Our results show that $\mathbf{h}_{11.10}$ operates based on several induction heads (Elhage et al. 2021) ($\mathbf{h}_{5.8}$, $\mathbf{h}_{6.1}$, $\mathbf{h}_{6.15}$ and $\mathbf{h}_{7.2}$) formed at $[m_{2}]$. These heads attend to the $[[m_{1}]+1]$ token due to the $m_{1}\equiv m_{2}$ conditioned matching operation, likely containing $m_{1}$-related information from previous token heads. We conclude, therefore, that $m_{1}$ and $m_{2}$ information are aggregated at position $[p]$. Additional details of the path patching results are available in the Appendix F.
Suppressive mechanism is revealed through Logit Lens.
To better understand the internal mechanism occurring at the $[p]$ position, we investigate attention head $\mathbf{h}_{11.10}$. Having previously examined the attention pattern (composed of attention weights), we now focus on the attention value and output by analyzing the OV circuit ($W_{V}W_{O}$(Elhage et al. 2021) using the logit lens method (Nostalgebraist 2020). Interestingly, the result in Figure 4(c) shows a clear negative diagonal pattern suggesting that $\mathbf{h}_{11.10}$ strongly suppresses the logit when attending to the same token as the corresponding output. Given our previous findings that $\mathbf{h}_{11.10}$ reads information from the subspace of token $[m_{2}]$, we conclude that it applies a suppressive mechanism to $m$-related information and writes it back to the residual stream’s subspace at $[p]$. If later heads carry this information to the last token, this mechanism becomes crucial for the model to arrive at the correct answer. For simplicity, we name head $\mathbf{h}_{11.10}$ as $m$-suppression head.
Localization of Term-Related Information Flow
Key information is moved from term-specific positions to the last position.
Now, we use the all-term intervention method to localize mover heads that carry term-related information. In the residual stream patching results (Figure 4(d)), we observe the highest positive score at the $[p]$ position, while negative scores are most prominent at the $[m_{1}]$ and $[m_{2}]$ positions, indicating that the information residing in these positions indeed contributes to their corresponding token prediction. This observation aligns with the importance of token embedding information at their respective positions given iterative refinement on the residual stream (Simoulin and Crabbé 2021). A closer examination of the last token position reveals that the negative effect propagates from relatively early layers (approximately layer 10 onwards), yet positive effects are from later layers (approximately layer 14 onwards), incurring a positive shift of the model’s prediction from $m$ to $p$.
Information is carried by later positive and negative mover heads.
Given the findings that the information for prediction resides in term-specific positions, we trace which attention heads transfer the information from each term position ($[p]$, $[m_{1}]$ and $[m_{2}]$) to the last token. We call these heads mover heads as the existence of “positive or negative mover heads” that carry or suppress information from the specific token position to the last token position (Wang et al. 2023; García-Carrasco, Maté, and Carlos Trujillo 2024). To discover the source of information flow, we investigate the heads with high attention value weight ($W_{V}$) patching score and identify 9 notable mover heads (see Appendix G for details of mover heads localization). By further investigation of attention patterns (Figure 4(f)–(h)), $\mathbf{h}_{14.14},\mathbf{h}_{15.14}$ and $\mathbf{h}_{18.12}$ are classified as positive copy heads attending highly to $[p]$, and $\mathbf{h}_{19.1}$ is a positive suppression head by attending to $[m_{2}]$ and $[s]$. $\mathbf{h}_{9.9}$, $\mathbf{h}_{11.1}$, $\mathbf{h}_{12.1}$, $\mathbf{h}_{17.2}$ and $\mathbf{h}_{23.10}$ are negative copy heads giving a strong attention to non-$[p]$.
A long induction mechanism in early layer movers causes biases towards the middle term.
[…]
Summary
Overall, we found that the internal mechanisms to arrive at the correct conclusion involve the following phases:
  1. 1.
    Long Induction: Early layers exhibit biases towards the wrong conclusion due to long-range repetition of the first premise “All A are B”.
  2. 2.
    Duplication: Induction heads aggregate information about duplicated middle terms in the premises.
  3. 3.
    Suppression: The model aggregates and inhibits middle-term information (i.e.,“B”) suppressing the long induction mechanism.
  4. 4.
    Mover: Token-specific information is propagated to the last token position. The process culminates in the prediction shift from B’ to the correct token ‘C’.
Therefore, the discovered circuit is mainly characterized by an internal error correction mechanism. Interestingly, this mechanism is different from the way humans would reason on syllogistic arguments through the use of abstract logical primitives and inference rules.
Circuit Evaluation
To evaluate the comprehensive correctness of the circuit, we assess necessity and sufficiency via the ablation method described in section 2. The ablation study in Figure 5(a) shows that the identified circuit is both necessary and sufficient, demonstrating a consistent performance degradation when removing circuit components and revealing a complete restoration of the original model’s performance when considering only the circuit’s subcomponents.
Figure 5: Circuit ablation results for (a) correctness and (b) robustness on the symbolic dataset. Solid lines show numeric perturbation-based performance, dotted lines indicate quantifier perturbation-based performance in (b), and the dashed line shows the baseline logit difference without knockouts.
Figure 6: Circuit ablation results for evaluating transferability on non-symbolic datasets: the belief-consistent (a) and the belief-inconsistent (b). The dashed line represents the baseline logit difference without knockouts.
We further verify the robustness of the symbolic circuit to superficial and semantic-preserving perturbations. In particular, we modify the letters into numbers (Figure 5(b)) and adopt semantically equivalent quantifiers and related verbs (e.g., “All … are” is converted into “Each … is”). The ablation result demonstrates that both types of perturbations do not undermine the sufficiency and necessity property of the circuit.
Circuit Transferability
The circuit is necessary for non-symbolic syllogistic arguments, yet not sufficient for belief-inconsistent ones.
We present ablation results for the two generated non-symbolic datasets: belief-consistent (Figure 6(a)) and belief-inconsistent (Figure 6(b)) (See Appendix H for the intervention results). Notably, the symbolic circuit proves necessary for both types of non-symbolic inputs, suggesting that the logic derived from symbolic syllogisms remains essential even when natural words are substituted. Regarding sufficiency, while belief-consistent data show significant performance recovery, we observe an inability to restore original performance on the belief-inconsistent set despite an increasing trend. These results indicate that belief biases encoded in different attention heads may play an important role.
Belief biases corrupt reasoning mechanisms.
To further investigate the impact of belief biases, we again conduct intervention experiments. Specifically, we observe that in the AAA-1 syllogism schema $(s,m_{1},m_{2},p)$, the subject token $s$ should be irrelevant for deriving the correct answer $p$ over $m$. Therefore, we leverage this property to verify whether word-specific biases are introduced when instantiating the schema with concrete knowledge. To this end, we perform Activation Patching by corrupting the subject term, transforming $(s,m_{1},m_{2},p)$ to $(s^{\prime},m_{1},m_{2},p)$, and measuring the degradation in logit difference as $\delta_{+}(p,m)-\delta_{-}(p,m)$. Notably, the non-symbolic setting exhibits a significant performance degradation of $0.66\pm 1.27$, representing a 299.96% change from the baseline. In contrast, the symbolic setting shows a minimal degradation of $0.00\pm 0.64$, a mere 0.35% drop. This drastic difference supports our hypothesis that the knowledge acquired during pre-training corrupts the content-independent reasoning circuit identified on the symbolic set with additional attention heads.
Generalization to Different Syllogistic Schemes
Mood-Figure C1 C2 C3 Acc
AII-3 (Datisi) $\checkmark$ $\checkmark$ $\checkmark$ 0.84
IAI-3 (Disamis) $\checkmark$ $\checkmark$ $\checkmark$ 0.68
IAI-4 (Dimaris) $\checkmark$ $\checkmark$ $\checkmark$ 0.68
AAA-1 (Barbara) $\checkmark$ $\checkmark$ $\checkmark$ 0.67
EAE-1 (Celarent) $\checkmark$ - $\checkmark$ 0.59
EIO-4 (Fresison) $\checkmark$ - $\checkmark$ 0.53
EIO-3 (Ferison) $\checkmark$ - $\checkmark$ 0.53
AII-1 (Darii) $\checkmark$ $\checkmark$ - 0.43
AOO-2 (Baroco) - - - 0.24
AEE-4 (Camenes) - - - 0.24
OAO-3 (Bocardo) $\checkmark$ - - 0.22
EIO-1 (Ferio) $\checkmark$ - - 0.18
EIO-2 (Festino) - - - 0.09
EAE-2 (Cesare) - - - 0.08
AEE-2 (Camestres) - - - 0.04
Table 2: Generalizability across unconditionally valid syllogistic forms. Columns C1, C2, and C3 indicate whether conditions are met ($\checkmark$) or not (-): C1: Necessity, C2: Sufficiency, C3: Positive Logit Difference. ‘Accuracy (Acc)’ shows performance on the completion task for each syllogism.
We further aim to understand whether the symbolic circuit is specific to the AAA-1 syllogism (Barbara). To this end, we extend our experiment to encompass all 15 unconditionally valid syllogisms (see Appendix A and I for details about the syllogistic schemes and more experimental results). To evaluate the transferability to all 15 schemes, we verify three main conditions, reported in Table 2: (C1) Necessity, (2) Sufficiency, and (C3) Positive Logit Difference. We also measure the accuracy of the completion task for each syllogism. From Table 2, we observe that the circuit is sufficient and necessary for all the syllogistic schemes in which the model achieves high downstream accuracy ($\geq$ 60%), supporting the conclusion that the circuit includes components that are crucial for the emergence of syllogistic reasoning in general. Notably, all three conditions are satisfied for the affirmative syllogisms (AII-3, IAI-3, IAI-4), for which the model achieves an accuracy above 0.6.
Generalization to Different Model Sizes
Finally, we expand our analysis to different sizes of GPT-2 (small, large and XL), assuming core mechanisms would transfer across models trained on the same dataset. Here, the intervention results show similar suppression mechanism patterns and information flow across all models. However, the residual stream of GPT-2 small in the all-term corruption setup is reversed due to its low downstream accuracy (see Appendix J). Moreover, as the model size increases, we found evidence that the contribution of attention heads becomes more complex. We hypothesize this might be caused by a stronger impact of world knowledge with increasing size, as suggested by the decrease in accuracy on the symbolic dataset and the increase on the non-symbolic set observed for the XL model (See Figure 3).
Related Work
Mechanistic circuit analysis has emerged as a promising approach to interpreting the internal mechanisms of Transformers (Olah et al. 2020; Nanda et al. 2023; Olsson et al. 2022; Wang et al. 2023; García-Carrasco, Maté, and Carlos Trujillo 2024; Hanna, Liu, and Variengien 2024). Existing approaches investigating internal reasoning mechanisms mainly focus on math-related tasks, elucidating the information flow for answering mathematical questions (Stolfo, Belinkov, and Sachan 2023) and examining arithmetic operations (Quirke and Barez 2024). Most pertinent to our work, Wiegreffe et al. (2024) provides a mechanistic interpretation of multiple-choice question answering, investigating attention head-level patterns. In general, our mechanistic analysis complements recent work investigating the challenges in processing reasoning arguments that contradict established beliefs and whether the reasoning in LMs stems from internalized rules or memorized content (Ando et al. 2023; Yu, Merullo, and Pavlick 2023; Wu et al. 2024; Talmor et al. 2020; Wu et al. 2024; Kassner, Krojer, and Schütze 2020; Haviv et al. 2023; Feldman 2020; Monea et al. 2024; Yu, Merullo, and Pavlick 2023; Wallat, Singh, and Anand 2020). To address this question, different mechanistic approaches have been adopted to localizing factual associations (Meng et al. 2022; Geva et al. 2023; Dai et al. 2022) or assessing conditions for generalisation (Wang et al. 2024). However, to the best of our knowledge, we are the first to investigate content-independent mechanisms for syllogistic reasoning.
Conclusion
In this study, we have provided a comprehensive mechanistic interpretation of syllogistic reasoning within auto-regressive language models. Overall, our findings suggest that LMs indeed learn transferable content-independent reasoning mechanisms but that, at the same time, such mechanisms might be contaminated and suppressed by the same specific world knowledge acquired during pre-training.
It is important to acknowledge some of the limitations in our study: (1) our analysis is conducted predominantly on the transitive property and term-specific information and does not consider the full spectrum of the reasoning dynamics that might appear in more complex scenarios. (2) Our experimentation is confined to a specific reasoning template that lacks the complexity and noise of real-world problems due to the intricate level of granularity and the computational complexity required in circuit analysis. Despite these limitations, we believe our findings could offer valuable insights into the reasoning mechanisms adopted by auto-regressive language models and lay a solid foundation for future research in this field.
Acknowledgements
This work was partially funded by the Swiss National Science Foundation (SNSF) project NeuMath: Neural Discourse Inference over Mathematical Texts (200021_204617).
Appendix
This document serves as a technical appendix to the paper titled A Mechanistic Interpretation of Syllogistic Reasoning in Auto-Regressive Language Models. It provides additional theoretical background, experimental details, and supplementary results that support the main findings presented in the original paper.
Appendix A A. Unconditionally Valid Syllogism Schemes
We list all 15 unconditionally valid syllogisms (Table 3).
Name Mood Figure Premise (m) Premise (M) Conclusion
Barbara AAA 1 $\forall x(s,m)$ $\forall x(m,p)$ $\forall x(s,p)$
All A are B. All B are C. All A are C.
Celarent EAE 1 $\forall x(s,m)$ $\forall x\neg(m,p)$ $\forall x\neg(s,p)$
All A are B. No B are C. No A are C.
Darii AII 1 $\exists x(s,m)$ $\forall x(m,p)$ $\exists x(s,p)$
Some A are B. All B are C. Some A are C.
Ferio EIO 1 $\exists x(s,m)$ $\forall x\neg(m,p)$ $\exists x\neg(s,p)$
Some A are B. No B are C. Some A are not C.
Camestres AEE 2 $\forall x\neg(s,m)$ $\forall x(p,m)$ $\forall x\neg(s,p)$
No A are B. All C are B. No A are C.
Cesare EAE 2 $\forall x(s,m)$ $\forall x\neg(p,m)$ $\forall x\neg(s,p)$
All A are B. No C are B. No A are C.
Baroco AOO 2 $\exists x\neg(s,m)$ $\forall x(p,m)$ $\exists x\neg(s,p)$
Some A are not B. All C are B. Some A are not C.
Festino EIO 2 $\exists x(s,m)$ $\forall x\neg(p,m)$ $\exists x\neg(s,p)$
Some A are B. No C are B. Some A are not C.
Disamis IAI 3 $\forall x(m,s)$ $\exists x(m,p)$ $\exists x(s,p)$
All B are A. Some B are C. Some A are C.
Datisi AII 3 $\exists x(m,s)$ $\forall x(m,p)$ $\exists x(s,p)$
Some B are A. All B are C. Some A are C.
Ferison EIO 3 $\exists x(m,s)$ $\forall x\neg(m,p)$ $\exists x\neg(s,p)$
Some B are A. No B are C. Some A are not C.
Bokardo OAO 3 $\forall x(m,s)$ $\exists x\neg(m,p)$ $\exists x\neg(s,p)$
All B are A. Some B are not C. Some A are not C.
Dimaris IAI 4 $\forall x(m,s)$ $\exists x(p,m)$ $\exists x(s,p)$
All B are A. Some C are B. Some A are C.
Camenes AEE 4 $\forall x\neg(m,s)$ $\forall x(p,m)$ $\forall x\neg(s,p)$
No B are A. All C are B. No A are C.
Fresison EIO 4 $\exists x(m,s)$ $\forall x\neg(p,m)$ $\exists x\neg(s,p)$
Some B are A. No C are B. Some A are not C.
Table 3: 15 Unconditionally valid syllogisms schemes. The table lists the syllogisms by their traditional names, moods, and figures, with formalised logical expressions on the first line and corresponding natural language examples on the second line. The minor premise (m) is presented before the major premise (M) to emphasise the transitive property inherent in these forms.
Appendix B B. Dataset Generation
B.1. Symbolic
The symbolic dataset comprises sentences where all terms in the premises are represented by abstract symbols (uppercase alphabet letters). From the set of all 26 uppercase alphabets, three-letter triples (e.g., A, B and C) are randomly sampled. For each triple, six permutated prompts and label pairs are generated following a template designed to minimise latent semantic interference among alphabet symbol tokens.
B.2. Non-symbolic
The non-symbolic datasets are constructed based on GenericsKB (Bhakthavatsalam, Anastasiades, and Clark 2020), a resource that provides a foundation for evaluation of sentence veracity with associated truthfulness scores (0-1). The dataset construction process involves the following steps:
  • Extraction: Select generic sentences with a truthfulness score of 1 based on GenericsKB (Bhakthavatsalam, Anastasiades, and Clark 2020), specifically those in the form A are B, using regular expression.
  • Syllogism Construction: We form universal affirmative syllogism arguments (Barbara) based on a template by chaining sentences where the predicate of one sentence logically matches the subject of another.
  • Constraints: We exclude terms tokenised into multiple tokens to maintain consistency in comparison with the symbolic dataset, which is essential for coherent circuit analysis.
  • Classification: To address issues of partial inclusion and syntactic ambiguity in constructed syllogism arguments, we employ GPT-4 (OpenAI et al. 2024) to classify whether arguments contain only truthful premises and whether the middle-terms are syntactically and semantically equivalent. This step helps us classify a belief-consistent dataset and a belief-inconsistent dataset.
  • Validation: We manually evaluate the alignment of classified arguments with human belief-consistency (i.e., premises are true and logic is valid).
This process ensures that our non-symbolic dataset maintains logical equivalence with the symbolic dataset while incorporating meaningful semantic real-word concepts.
B.3. Data Statistics
We organise data statistics for generated datasets (Table 4).
Statistic SYM BC BI
Number of Samples 90 90 90
Token Sequence Length 15 15 15
Unique Terms ($s$) 26 87 86
Unique Terms ($m$) 26 70 35
Unique Terms ($p$) 26 59 64
Table 4: Summary of Dataset Statistics. SYM refers to the symbolic dataset, BC refers to the non-symbolic belief-consistent dataset, and BI refers to the non-symbolic belief-inconsistent dataset.
Appendix C C. Embedding Space Analysis Details
One well-known method for analysing hidden activation in transformer language models is projecting activation into embedding space, the so-called logit lens (Nostalgebraist 2020; Elhage et al. 2021; Geva et al. 2022). In particular, we focus on an input-agnostic approach (Dar et al. 2023; Hanna, Liu, and Variengien 2024), utilising both unembedding ($\mathbf{W}_{U}$) and embedding ($\mathbf{W}_{E}$) matrices to construct a $\mathbb{R}^{|V|\times|V|}$ matrix, where $|V|$ represents the model’s vocabulary size. In this study, we employ the OV circuit (Elhage et al. 2021) of attention heads, formed by attention value and output weights ($W_{V}W_{O}$), to understand the general behaviour of how source tokens influence output logits. The OV circuit-based logit lens for attention head $\mathbf{h}$ is formalised as $\mathbf{W}_{E}\mathbf{W}^{\mathbf{h}}_{V}\mathbf{W}_{O}^{\mathbf{h}}\mathbf{W}_{U}\in\mathbb{R}^{|V|\times|V|}$. For simplicity and in line with established conventions (Elhage et al. 2021; Dar et al. 2023), this formulation ignores layer normalisation.
Appendix D D. Circuit Ablation Method Details
To measure the necessity and sufficiency of the circuit $\mathcal{C}$ in the model $\mathcal{M}$, we knock out attention heads in $\mathcal{H}$ from the model $\mathcal{M}$ and measure the average logit difference $\delta(p,m)$ along the batch. We denote the logit difference in circuit state $\mathcal{C}$ as $\delta(p,m,\mathcal{C})$ and every subset of heads set as $H\subset\mathcal{H}$.
In order to verify the head’s necessity in the model, we conduct a cumulative ablation of $\mathcal{C}$ from total circuit $\mathcal{M}$, progressing from downstream to upstream layers. At each ablation step $k$, where $\mathcal{C}_{k}^{*}=\mathcal{M}\setminus H_{k}$ and $H_{k}$ denotes the set of ablated attention heads up to step $k$ as:
$\mathbb{E}_{x\sim X}[\delta_{k}(p,m,\mathcal{C}_{k}^{*})]\quad\text{where}\quad\mathcal{C}_{k}^{*}=\mathcal{M}\setminus H_{k}$
Conversely, we perform a cumulative addition of attention heads for evaluating the sufficiency of the circuit, starting from earlier layers and progressing to later ones, while maintaining all other attention heads in a mean-ablated state. At each addition step $j$, where $\mathcal{C}_{j}^{*}=\mathcal{M}\setminus(\mathcal{H}\setminus H_{j})$ and $H_{j}$ denotes the set of added attention heads up to step $j$ as:
$\mathbb{E}_{x\sim X}[\delta_{j}(p,m,\mathcal{C}_{j}^{*})]\quad\text{where}\quad\mathcal{C}_{j}^{*}=\mathcal{M}\setminus(\mathcal{H}\setminus H_{j})$
Appendix E E. GPT-2 Model Architecture
We provide a self-contained concise overview of the GPT-2 model architecture, highlighting the main components and their mathematical relationships. Bias terms are not presented for the simplicity. We refer the conventions of notation in Elhage et al. (2021) and (Geva et al. 2023).
Notation and Preliminaries
  • $x_{i}$ One-hot encoded vector representing the $i$-th token in the input sequence.
  • $W_{E}^{\text{token}},W_{E}^{\text{pos}}$ Token and positional embedding matrices.
  • $W_{Q},W_{K},W_{V}$ Query, key, and value weight matrices of attention heads.
  • $W^{\text{in}},W^{\text{out}}$ Input and output weight matrices of MLP feed-forward networks.
  • $d$ Dimension of the attention head state embedding vectors.
  • $D$ Dimension of the hidden state embedding vectors.
  • $N$ Length of the input sequence.
  • $L$ Total number of transformer layers.
  • $\mathcal{V}$ Vocabulary set of the model.
  • $\sigma$ and $\sigma^{\prime}$ Activation functions used in self-attention and MLP sublayers, respectively.
Embedding Layer.
Each token in the input sequence is transformed into an embedded representation by combining token-specific and positional embeddings:
$r_{i}^{0}=x_{i}W_{E}^{\text{token}}+x_{i}W_{E}^{\text{pos}}$
This operation initializes the input for the transformer layers, where $r_{i}^{0}$ represents the initial embedded state of the $i$-th token.
Residual Stream.
The residual stream facilitates the propagation of information across transformer layers. It is updated at each layer by contributions from the multi-head self-attention and multi-layer perceptron sublayers:
$r_{i}^{l}=r_{i}^{l-1}+A_{i}^{l}+M_{i}^{l}$
where $r_{i}^{l}$ represents the state of the $i$-th token at layer $l$, and $A_{i}^{l}$, $M_{i}^{l}$ are outputs from the Multi-Head Self-Attention (MHSA) and Multi-Layer Perceptron (MLP) sublayers.
Multi-Head Self-Attention (MHSA).
This sublayer processes information by applying self-attention mechanisms across multiple ‘heads’ of attention, enabling the model to capture various aspects of the input data, where $\sigma$ represent the non-linearity function:
$A_{i}^{l}=\sum_{h\in H}\left(\sigma\left(\frac{(r_{i}^{l,h}W^{h}_{Q})(r_{i}^{l,h}W^{h}_{K})^{T}}{\sqrt{d_{K}}}\right)(r_{i}^{l,h}W^{h}_{V})\right)W^{h}_{O}$
Each head computes a weighted sum of all tokens’ transformed states, focusing on different subsets of sequence information.
Multi-Layer Perceptron (MLP).
Each token’s representation is further processed in a position-wise manner by the MLP sublayer:
$M_{i}^{l}=\sigma^{\prime}(r_{i}^{l}W^{in})(W^{out})^{T}$
The MLP modifies each token’s state locally, enhancing its ability to process information.
Layer Normalisation.
Before processing by MHSA and MLP, each token’s state is normalised to stabilise learning and reduce training time:
$\text{LN}(r_{i}^{l})=\gamma\odot\frac{r_{i}^{l}-\mu_{i}^{l}}{\sqrt{(\sigma_{i}^{l})^{2}+\epsilon}}+\beta$
This step ensures that the activations across different network layers maintain a consistent scale.
Prediction Head.
The prediction head generates logits for the next token prediction using the final transformed states:
$\text{logits}_{N}=\text{LN}(r_{N}^{L})W_{U}$
The predicted token is chosen based on the sampling method (e.g., greedy decoding) at the last position $N$.
Appendix F F. Path Patching Details
Path patching, an generalised version of activation patching, aims to compute direct effects among model components, rather than indirect effects diffused from intervened components (Wang et al. 2023). This method involves freezing non-targeted activations during an initial forward pass, storing the corrupted targeted activation state, and then executing a subsequent forward pass with the targeted activation state substituted by the corrupted one. This approach isolates targeted component-to-component effects, eliminating non-relevant interactions. In our implementation, we employ a noising method where clean activations are replaced with corrupted ones, resulting in negative scores indicating positive contributions from corresponding components. Our findings, illustrated in Figure 7, reveal that $\mathbf{h}_{11.10}$ strongly operates based on several heads: $\mathbf{h}_{5.8}$, $\mathbf{h}_{6.1}$, $\mathbf{h}_{6.15}$, and $\mathbf{h}_{7.2}$. Manual investigation of attention patterns subsequently confirmed these as induction heads (Elhage et al. 2021).
Figure 7: Path patching results for (Top) symbolic inputs and non-symbolic inputs (Bottom) in the middle-term corruption setup. Note that negative heads are positively influencing heads, as we replace corrupted activation with clean ones in our noising approach.
Appendix G G. Localising Mover Heads
To efficiently classify the numerous attention heads (384 in GPT-2 medium) within the model, we calculate the Positional Patching Difference (PPD) to construct four-quadrant analysis. This method involves intervening in the attention value weights $W_{V}$ and constructs a distribution layout of heads, providing a systematic classification for quadrant groups. The layout is composed of the PPD score ($|S[p]|-|S[m_{1}]+S[m_{2}]|$) on the x-axis and the patching score for all sequence positions ($\mathcal{S}$) on the y-axis.
Based on this PPD score quadrant layout analysis, we categorise heads according to their quadrant positions, revealing insights into their functional roles:
  • (1)
    First quadrant (Positive Copy Candidates): Heads with $\mathcal{S}>0$ and positive PPD, predominantly copying $[p]$-based information.
  • (2)
    Second quadrant (Positive Suppression Candidates): Heads with $\mathcal{S}>0$ but negative PPD, predominantly suppressing [$m$]-based information.
  • (3)
    Third quadrant (Negative Copy Candidates): Heads with $\mathcal{S}<0$ and negative PPD, predominantly copying [$m$]-based information.
  • (4)
    Fourth quadrant (Negative Suppression Candidates): Heads with $\mathcal{S}<0$ but positive PPD, predominantly suppressing $[p]$-based information.
The intervention of attention head values ($W_{V}$) is crucial because all-term corruption maintains the same positional alignment for all samples $(s,m_{1},m_{2},p)\rightarrow(s^{\prime},m_{1}^{\prime},m_{2}^{\prime},p^{\prime})$, resulting in minimal impact from attention head pattern interventions composed by attention query and key. Consequently, we can infer that the attention head output patching results primarily derive from the attention value activation. This property enables us to localise the source token positions from which the term information is moved by attention value patching.
Formally, we denote the attention query, key, value, and output matrix activations as $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, $\mathbf{W}_{V}$, and $\mathbf{W}_{O}$ respectively. One attention head output can be expressed as $\sigma\left((\mathbf{R}\mathbf{W}_{Q})(\mathbf{R}\mathbf{W}_{K})^{T}\right)(\mathbf{R}\mathbf{W}_{V})\mathbf{W}_{O}$, where $\sigma$ represents the non-linearity function (softmax in GPT-2) applied in the attention pattern and $\mathbf{R}$ denotes the residual stream.
To include only influential heads in the circuit, we extracted outlier heads with an absolute patching score exceeding the threshold ($\tau$), defined by the mean and standard deviation of all heads’ scores as:
$\{\mathbf{h}\mid|\mathcal{S}_{\mathbf{h}}|>\tau\}\quad\text{where}\quad\tau=\mu+2\sigma$
Our analysis identifies 9 heads for the symbolic circuit, as illustrated by red dots in Figure 8(a). $\mathbf{h}_{14.14}$, $\mathbf{h}_{15.14}$, and $\mathbf{h}_{18.12}$ are classified as positive copy heads, while $\mathbf{h}_{19.1}$ is grouped as a positive suppression head. $\mathbf{h}_{9.9}$, $\mathbf{h}_{11.1}$, $\mathbf{h}_{12.1}$, $\mathbf{h}_{17.2}$, and $\mathbf{h}_{23.10}$ are classified as negative copy heads. It is noteworthy that non-symbolic results reveal a more complex and noisy pattern in the mover localisation process (Figure 8(b)).
Figure 8: Distribution of attention heads in PPD-based quadrant analysis. (a) Results for symbolic inputs. (b) Results for non-symbolic inputs. The x-axis represents normalised PPD scores, while the y-axis represents normalised attention value-based patching scores ($S$). Red dots indicate attention heads above the threshold ($\tau$) of patching score.
Appendix H H. Non-symbolic Circuit Analysis
We present the results of our non-symbolic circuit analysis (Figure 9).
Figure 9: Comprehensive results of the non-symbolic circuit analysis. (a) Attention output patching results and (b) residual stream patching results in the middle-term intervention setup. (c) OV circuit logit lens results for $\mathbf{h}_{11.10}$, with input and output comprising 26 uppercase letters. (d) Residual stream patching results in the all-term corruption setup. (e)-(h) Average attention patching weights across the batch: (e) weights for token position $[p]$ attending, and (f)-(h) weights for the last token position attending. For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.
Appendix I I. Generalisability to Other Syllogisms
To assess the generalisability of our model, we applied the circuit ablation method to all 15 unconditionally valid syllogisms (Table 3) and the result is shown in Figure 10. Unconditionally valid syllogisms are logical arguments that maintain their validity irrespective of the truth values of their premises. The validity of these syllogisms is determined solely by their logical form, which is characterised by mood and figure combinations. The mood of a syllogism is defined by the arrangement of four proposition types (”All (A)”, ”No (E)”, ”Some (I)”, ”Some … not (O)”) across its two premises and conclusion. Conversely, the figure of a categorical syllogism is determined by the structural arrangement of terms within the constituent sentences. This approach enables the evaluation of syntactic logic in isolation from contextual interference and belief-bias effects.
Figure 10: Necessity and sufficiency performance results from circuit ablation method for all unconditionally valid syllogistic forms. Labels (e.g., AII-3) denote mood and figure combinations.
Appendix J J. Generalisability Across Model Sizes
To assess the generalisability of our findings across different model capacities, we extended our symbolic circuit analysis to various GPT-2 model sizes. Figure 11 presents the results of this comparative analysis.
Figure 11: Comprehensive results of the symbolic circuit analysis across different model sizes. (a) Attention output patching results and (b) residual stream patching results in the middle-term intervention setup. (c) OV circuit logit lens results for $m$-suppression head, with input and output comprising 26 uppercase letters. (d) Residual stream patching results in the all-term corruption setup. For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.

### v2

@@
 Abstract
-Recent studies on logical reasoning in auto-regressive Language Models (LMs) have sparked a debate on whether such models can learn systematic reasoning principles during pre-training or merely exploit superficial patterns in the training data.
-This paper presents a mechanistic interpretation of syllogistic reasoning in LMs to further enhance our understanding of internal dynamics.
-Specifically, we present a methodology for circuit discovery aimed at disentangling content-independent reasoning mechanisms from world knowledge acquired during pre-training.
+Recent studies on logical reasoning in Language Models (LMs) have sparked a debate on whether they can learn systematic reasoning principles during pre-training or merely exploit superficial patterns in the training data.
+This paper presents a mechanistic interpretation of syllogistic reasoning in LMs to advance the understanding of internal dynamics.
+Specifically, we present a methodology for circuit discovery aimed at interpreting content-independent reasoning mechanisms.
 Through two distinct intervention methods, we uncover a sufficient and necessary circuit involving middle-term suppression that elucidates how LMs transfer information to derive valid conclusions from premises.
 Furthermore, we investigate how belief biases manifest in syllogistic reasoning, finding evidence of partial contamination from additional attention heads responsible for encoding commonsense and contextualized knowledge.
-Finally, we explore the generalization of the discovered mechanisms across various syllogistic schemes and model sizes, finding that the identified circuit is sufficient and necessary for all the schemes on which the model achieves high downstream accuracy ($\geq$ 60%).
-Overall, our findings suggest that LMs indeed learn transferable content-independent reasoning mechanisms, but that, at the same time, such mechanisms do not involve generalisable and abstract logical primitives, being susceptible to contamination by the same world knowledge acquired during pre-training.
-Introduction
-Transformer-based Language Models (LMs) have led to remarkable results across various natural language processing tasks (Radford et al.
-2018, 2019; Brown et al.
-2020; OpenAI et al.
-2024; Jason et al.
-2022; Bubeck et al.
-2023).
-This success has catalyzed research interest in systematically exploring the reasoning capabilities emerging during pre-training (Clark, Tafjord, and Richardson 2020).
-Recent findings suggest that logical reasoning abilities may emerge in large-scale models (Rae et al.
-2022; Kojima et al.
-2022; Wei et al.
-2022) or through transfer learning on specialized datasets (Betz, Voigt, and Richardson 2021).
-However, ongoing debates question whether these models apply systematic inference rules or reuse superficial patterns learned during pre-training (Talmor et al.
-2020; Kassner, Krojer, and Schütze 2020; Wu et al.
-2024).
-This controversy underscores the need for a deeper understanding of the low-level logical inference mechanisms in LMs (Rozanova, Valentino, and Freitas 2024; Rozanova et al.
-2023, 2022).
+Finally, we explore the generalization of the discovered mechanisms across various syllogistic schemes, model sizes and architectures, finding that the identified circuit is sufficient and necessary for the schemes on which the models achieve high downstream accuracy ($\geq$ 60%), and that the activation patterns apply to models of different families.
+Overall, our findings suggest that LMs indeed learn transferable content-independent reasoning mechanisms, but that, at the same time, such mechanisms do not involve generalizable and abstract logical primitives, being susceptible to contamination by the same world knowledge acquired during pre-training.
+A Mechanistic Interpretation of Syllogistic Reasoning in
+Auto-Regressive Language Models
+Geonhee Kim1, Marco Valentino3, André Freitas1,2,3 1Department of Computer Science, University of Manchester, UK 2National Biomarker Centre (NBC), CRUK Manchester Institute, UK 3 Idiap Research Institute, Switzerland
+1 Introduction
 Figure 1: Conceptual representation of the circuit for processing symbolic syllogisms: (Long Induction) Early layers exhibit biases towards the wrong conclusion due to long-range repetition of the first premise “All A are B”.
@@
 The process culminates in the prediction shift from ‘B’ to the correct token, ‘C.’
-To improve our understanding of the internal mechanisms, this paper focuses on Mechanistic Interpretability (Olah et al.
-2020; Nanda et al.
-2023), aiming to discover the core circuit responsible for content-independent syllogistic reasoning.
+Transformer-based Language Models (LMs) have led to remarkable results across various natural language processing tasks (Radford et al., 2018, 2019; Brown et al., 2020; OpenAI et al., 2024; Jason et al., 2022; Bubeck et al., 2023).
+This success has catalyzed research interest in systematically exploring the reasoning capabilities emerging during pre-training (Clark et al., 2020).
+Recent findings suggest that logical reasoning abilities may emerge in large-scale models (Rae et al., 2022; Kojima et al., 2022; Wei et al., 2022) or through transfer learning on specialized datasets (Betz et al., 2021).
+However, ongoing debates question whether these models apply systematic inference rules or reuse superficial patterns learned during pre-training (Talmor et al., 2020; Kassner et al., 2020; Wu et al., 2024).
+This controversy underscores the need for a deeper understanding of the low-level logical inference mechanisms in LMs (Rozanova et al., 2024, 2023, 2022).
+To improve our understanding of the internal mechanisms, this paper focuses on Mechanistic Interpretability (Olah et al., 2020; Nanda et al., 2023), aiming to discover the core circuit responsible for syllogistic reasoning.
 In particular, we focus on categorical syllogisms with universal affirmative quantifiers (i.e., AAA-1, Barbara) motivated by two key factors.
-First, as observed in natural logic studies (MacCartney and Manning 2007), this form of syllogistic reasoning is prevalent in everyday language.
+First, as observed in natural logic studies (MacCartney and Manning, 2007), this form of syllogistic reasoning is prevalent in everyday language.
 Therefore, it is likely that LMs are exposed to such reasoning schema during pre-training.
-Second, AAA-1 is a form of unconditionally valid syllogism independent of the premises’ truth condition (Holyoak and Morrison 2005).
-This characteristic offers a deterministic and scalable task design, other than allowing us to investigate the disentanglement between reasoning and knowledge representation.
-Through the adoption of Activation Patching (Meng et al.
-2022) and embedding space analysis (i.e., Logit Lens) (Nostalgebraist 2020; Geva et al.
-2022; Dar et al.
-2023), we investigate the following main research questions: RQ1: What is the content-independent syllogistic reasoning mechanism internalized in LMs during pre-training?; RQ2: Are content-independent mechanisms disentangled from specific world knowledge and belief biases?; RQ3: Does the core reasoning mechanism generalize across model sizes and different syllogistic schemes?
+Second, AAA-1 is a form of unconditionally valid syllogism independent of the premises’ truth condition (Holyoak and Morrison, 2005).
+This characteristic offers a deterministic and scalable task design, other than allowing us to investigate the disentanglement between reasoning and knowledge representation Wysocka et al.
+(2024); Bertolazzi et al.
+(2024); Lampinen et al.
+(2024).
+Through mechanistic interpretability techniques such as Activation Patching (Meng et al., 2022) and embedding space analysis (i.e., Logit Lens) (Nostalgebraist, 2020; Geva et al., 2022; Dar et al., 2023), we investigate the following main research questions: RQ1: How is the content-independent syllogistic reasoning mechanism internalized in LMs during pre-training?; RQ2: Are content-independent mechanisms disentangled from specific world knowledge and belief biases?; RQ3: Does the core reasoning mechanism generalize across model sizes and syllogistic schemes?
 To answer these questions, we present a methodology that consists of 3 main stages.
@@
 We present the following overall conclusions:
-  1. 1.
-    The circuit analysis reveals that auto-regressive LMs develop specific inference mechanisms during pre-training, finding evidence supporting a three-stage mechanism for syllogistic reasoning: (1) naive recitation of the first premise; (2) suppression of duplicated middle-term information; and (3) mediation towards the correct output through the interplay of attention heads.
-  2. 2.
+1.
+The circuit analysis reveals that LMs develop specific inference mechanisms during pre-training, finding evidence supporting a three-stage mechanism for syllogistic reasoning: (1) naive recitation of the first premise; (2) suppression of duplicated middle-term information; and (3) mediation towards the correct output through the interplay of attention heads.
+2.
 Further experiments on circuit transferability demonstrate that the identified mechanism is still necessary for reasoning on syllogistic schemes instantiated with commonsense knowledge.
-However, a deeper analysis suggests that specific belief biases acquired during pre-training might play an important role, contaminating the content-independent circuit mechanism with additional attention heads responsible for encoding general and contextualized world knowledge.
-  3. 3.
+However, a deeper analysis suggests that specific belief biases acquired during pre-training might contaminate the content-independent circuit mechanism with additional attention heads responsible for encoding contextualized world knowledge.
+3.
 We found that the identified circuit is sufficient and necessary for all the syllogistic schemes in which the model achieves high downstream accuracy ($\geq$ 60%).
-This result suggests that LMs learn content-independent reasoning mechanisms transferable across abstract schemes.
-  4. 4.
-    The intervention results on models with different sizes show similar suppression mechanism patterns and information flow.
+This result suggests that LMs learn reasoning mechanisms that are transferable across different schemes.
+4.
+The intervention results on models with different architectures and sizes (i.e., GPT-2 (Radford et al., 2019), Pythia (Biderman et al., 2023), Llama (Grattafiori et al., 2024), Qwen (Yang et al., 2024)) show similar suppression mechanism patterns and information flow.
 However, we found evidence that the contribution of attention heads becomes more complex with model sizes, further supporting the hypothesis of increasing contamination from external world knowledge.
-Our following study is conducted using TransformerLens (Nanda and Bloom 2022) on an Nvidia A100 GPU with 80GB of memory.
-The dataset and code to reproduce our experiments are available online to encourage future work in the field [^1] http://anonynous˙url.com.
-Methodology
-Our main research objective is to discover and interpret the core mechanisms adopted by auto-regressive Language Models (LMs) when performing content-independent syllogistic reasoning.
-To this end, we present a methodology that consists of 3 main stages.
-First, we define a syllogistic completion task that can be instrumental for our analysis.
-Second, we leverage the syllogistic completion task to implement a circuit discovery pipeline on a syllogistic schema instantiated only with symbolic variables (Table 1, Symbolic).
-Third, we investigate the generalization of the identified circuit on concrete schemes instantiated with commonsense knowledge and explore how the internal behavior varies with different schemes and model sizes.
 Premises ($\mathcal{P}_{1},\mathcal{P}_{2}$) Conclusion ($\mathcal{C}$)
@@
 all pilots are blond.
-Table 1: Examples of the same syllogistic schema (i.e., AAA-1, Barbara) instantiated with different content (with the expected completion in bold).
-The ability to complete the syllogism is a function of the reasoning schema, independent of the specific instantiation or truth condition of the premises.
-Figure 2: The conceptual pipeline of the symbolic circuit analysis on the syllogism completion task.
-Figures (a)–(e) illustrate the stages involved in processing a syllogistic argument: argument formalization (a), construction of clean (b) and corrupted inputs (c), activation patching (d) and embedding space analysis (i.e., Logit Lens) (e).
-The uncovered symbolic circuit is subsequently evaluated to determine whether the core reasoning schema operates independently from world knowledge or belief bias.
-Syllogism Completion Task
-We design a syllogism completion task for assessing the reasoning abilities of LMs, building upon established approaches in the literature (Betz, Voigt, and Richardson 2021; Wu et al.
-2023).
+Table 1: Examples of a syllogistic schema (i.e., AAA-1, Barbara).
+The logical validity of a conclusion is only a function of the reasoning schema, being independent of the specific variables or truth condition of the premises.
+Figure 2: The conceptual pipeline of the circuit discovery methodology.
+Figures (a)–(e) illustrate the stages in processing a syllogistic argument: formalization (a), construction of clean (b) and corrupted inputs (c), activation patching (d) and embedding space analysis (i.e., Logit Lens) (e).
+The uncovered symbolic circuit is evaluated to determine whether the core reasoning schema operates independently from world knowledge or belief bias.
+2 Methodology
+Our main research objective is to discover and interpret the core mechanisms adopted by auto-regressive Language Models (LMs) when performing content-independent syllogistic reasoning.
+To this end, we present a methodology that consists of 3 main stages.
+First, we define a syllogistic completion task that can be instrumental for our analysis.
+Second, we leverage the syllogistic completion task to implement a circuit discovery pipeline on a syllogistic schema instantiated only with symbolic variables (Table 1, Symbolic).
+Third, we investigate the generalization of the identified circuit on concrete schemes instantiated with commonsense knowledge and explore how the internal behavior varies with different schemes and models.
+2.1 Syllogism Completion Task
+We design a syllogism completion task for assessing the reasoning abilities of LMs, building upon established approaches in the literature (Betz et al., 2021; Wu et al., 2023).
 In particular, we focus on categorical syllogisms with universal affirmative quantifiers (i.e., AAA-1, Barbara) because of their frequency in natural language and their content-independent reasoning property.
@@
 men and mortal), and $m_{1}$ and $m_{2}$ (e.g., humans) denoting the middle terms in the two premises, with $m_{1}\equiv m_{2}$ in the case of the AAA-1 syllogism.
-To evaluate syllogistic reasoning in LMs, we formalize the completion task as masking language modeling, removing the final token $p$ from the conclusion (e.g., All men are humans.
+To evaluate syllogistic reasoning in LMs, we formalize the completion task as language modeling, removing the final token $p$ from the conclusion (e.g., All men are humans.
 All humans are mortal.
@@
 On the other side, the non-symbolic dataset is constructed by replacing the letters with words while preserving the syllogistic schema.
-Here, we generate two different non-symbolic sets: a belief-consistent set containing true premises and a belief-inconsistent set containing false premises (see Table 1).
-To guarantee the truth condition of the premises, we leverage GenericsKB (Bhakthavatsalam, Anastasiades, and Clark 2020), a knowledge base containing statements about commonsense knowledge.
+Here, we generate two different non-symbolic sets: a belief-consistent set containing true premises and a belief-inconsistent set containing false premises to address RQ2 (see Table 1).
+To guarantee the truth condition of the premises, we leverage GenericsKB (Bhakthavatsalam et al., 2020), a knowledge base containing statements about commonsense knowledge.
 The detailed generation process is in Appendix B.
-Circuit Discovery
+2.2 Circuit Discovery
 Our main objective is to find a circuit $\mathcal{C}$ for syllogistic reasoning in a Language Model $\mathcal{M}$.
 A circuit $\mathcal{C}$ can be defined as a subset of the original model $\mathcal{M}$ that is both sufficient and necessary for achieving the original model performance on the syllogistic completion task.
-In order to identify a circuit, we employ Activation Patching together with circuit ablation methods (Meng et al.
-2022; Vig et al.
-2020).
+In order to identify a circuit, we employ Activation Patching together with circuit ablation (Meng et al., 2022; Vig et al., 2020).
 Activation Patching.
@@
 Therefore, all $s$ are [mask]”, and its correct completion $y=p$, the Activation Patching technique consists of the following steps:
-  1. (1)
-    Clean Run: For the target input $(s,m_{1},m_{2},p)$, we record the baseline logit difference, $\delta_{+}(p,m)$ from the forward pass output of the model (Figure 2(1)).
-  2. (2)
-    Corrupted Run: We re-run the model on a corrupted input after applying a specific intervention (e.g., changing the middle term $m_{2}$) and record the logit difference $\delta_{-}(p,m)$ (Figure 2(3)).
-  3. (3)
-    Intervened Run: We run the model with the corrupted inputs again and replace the corrupted activations with those from the clean runs to compute the response from the remaining components and measure the causal impact of the intervention.
+(1) Clean Run.
+For the target input $(s,m_{1},m_{2},p)$, we record the baseline logit difference, $\delta_{+}(p,m)$ from the forward pass output of the model (Figure 2(1)).
+(2) Corrupted Run.
+We re-run the model on a corrupted input after applying a specific intervention (e.g., changing the middle term $m_{2}$) and record the logit difference $\delta_{-}(p,m)$ (Figure 2(3)).
+(3) Intervened Run.
+We run the model with the corrupted inputs again and replace the corrupted activations with those from the clean runs to compute the response from the remaining components and measure the causal impact of the intervention.
 Here, we measure the adjusted logit difference $\delta_{p}(p,m)$ (Figure 2(2)).
-  4. (4)
-    Quantification: We quantify the causal impact of each intervention using a patching score $\mathcal{S}$ following (Heimersheim and Nanda 2024) as shown in Figure 2(4).
-This score is further normalized to $[-1,1]$.
-We complement Activation Patching with known methods for analysing hidden activations in transformers, including Logit Lens (Nostalgebraist 2020) with an input-agnostic approach (Dar et al.
-2023; Hanna, Liu, and Variengien 2024).
+(4) Quantification.
+We quantify the causal impact of each intervention using a patching score $\mathcal{S}$ following (Heimersheim and Nanda, 2024) as shown in Figure 2(4). This score is further normalized to $[-1,1]$.
+We complement Activation Patching with known methods for analysing hidden activations in transformers, including Logit Lens (Nostalgebraist, 2020) with an input-agnostic approach (Dar et al., 2023; Hanna et al., 2024).
 Additional technical details can be found in Appendix C.
-Interventions.
+Figure 3: Results of the symbolic circuit analysis.
+(a) Attention output patching and (b) residual stream patching obtained via middle-term intervention.
+(c) OV circuit logit lens results for attention head $\mathbf{h}_{11.10}$, with input and output comprising 26 uppercase letters.
+(d) Residual stream patching obtained via all-term corruption.
+(e)-(h) Average attention weights across the batch: (e) weights for token position $[p]$, and (f)-(h) weights for the last token position.
+For clarity, the symbol ‘-’ on the y-axis denotes the averaged values for tokens appearing between terms.
+2.3 Interventions
 The choice of tokens to corrupt is a critical aspect of the experimental design, and it is essential to establish an appropriate type of intervention that aligns with the hypothesis being tested.
 In this study, we employ two distinct interventions to isolate the mechanisms related to the reasoning schema from the propagation of specific token information:
-  1. (1)
-    Middle-Term Corruption: To investigate the reasoning mechanism employed in the syllogism completion task, we primarily focus on the transitive property of the middle term, $(m_{1}\rightarrow m_{2})\wedge(m_{1}\equiv m_{2})$.
+(1) Middle-Term Corruption.
+To investigate the reasoning mechanism employed in the syllogism completion task, we primarily focus on the transitive property of the middle term, $(m_{1}\rightarrow m_{2})\wedge(m_{1}\equiv m_{2})$.
 We hypothesize that disrupting the transitive property will localize the component responsible for syllogistic reasoning.
 Therefore, our intervention method replaces the second middle term $m_{2}$ with an unseen symbol $m_{2}^{\prime}$, breaking the equality and effectively corrupting the validity of the reasoning, pivoting the correct answer towards $m$: $(s,m_{1},m_{2},p)\rightarrow(s,m_{1},m_{2}^{\prime},p)$.
-  2. (2)
-    All-Term Corruption: We examine how term-related information flows to the last position for the final prediction to identify potential mover heads.
-To this end, we replace the original terms $(s,m_{1},m_{2},p)$ with different terms $(s^{\prime},m_{1}^{\prime},m_{2}^{\prime},p^{\prime})$ while keeping the completion answers unchanged.
-We hypothesize that if the mover heads carry the $p$ related information over $m$ related information for a valid prediction, the logit difference will increase; otherwise, it will decrease.
-Figure 3: Accuracy on the syllogism completion task across all scales of GPT-2. We observe a phase change from small to medium models in overall performance.
-Circuit Ablation.
+(2) All-Term Corruption.
+We examine how term-related information flows to the last position for the final prediction to identify potential mover heads.
+To this end, we replace the original terms $(s,m_{1},m_{2},p)$ with different terms $(s^{\prime},m_{1}^{\prime},m_{2}^{\prime},p^{\prime})$ while keeping the answers unchanged.
+We hypothesize that if heads carry the information preferring $p$ over $m$ for the valid prediction, the logit difference will increase; otherwise, it will decrease.
+2.4 Circuit Ablation.
 Once we discover a relevant circuit $\mathcal{C}$, we evaluate its necessity and sufficiency using mean ablation.
@@
 Additional details are included in Appendix D.
-Empirical Evaluation
-Figure 4: Comprehensive results of the symbolic circuit analysis.
-(a) Attention output patching results and (b) residual stream patching results in the middle-term intervention setup.
-(c) OV circuit logit lens results for $\mathbf{h}_{11.10}$, with input and output comprising 26 uppercase letters.
-(d) Residual stream patching results in the all-term corruption setup.
-(e)-(h) Average attention patching weights across the batch: (e) weights for token position $[p]$ attending, and (f)-(h) weights for the last token position attending.
-For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.
-Model Selection.
+3 Empirical Evaluation
+Empirical Setup.
 We select the LM for the circuit analysis based on a trade-off between model size and performance.
-To this end, we measure the accuracy of GPT-2 (Radford et al.
-2019) with different sizes (117M, 345M, 762M, and 1.5B) on the syllogism completion task.
-Specifically, we aim to identify potential phase transition points where significant performance changes emerge (Jason et al.
-2022; Kaplan et al.
-2020).
-As shown in Figure 3, we observe a notable phase transition from small to medium sizes, where average performance increases by 139.97% in the symbolic setting and 119.21% in the non-symbolic setting.
-Given these results, we select GPT-2 Medium for circuit discovery and subsequently test the compatibility with different model sizes (See Appendix E for the model architecture).
-For all experiments, we use 90 samples each for symbolic and non-symbolic arguments to balance the depth of analysis with computational efficiency.
-Localization of Transitive Reasoning Mechanisms
+To this end, we measure the accuracy of GPT-2 (Radford et al., 2019) with different sizes (117M, 345M, 762M, and 1.5B) on the syllogism completion task.
+Specifically, we aim to identify potential phase transition points where significant performance changes emerge (Jason et al., 2022; Kaplan et al., 2020).
+As shown in Figure 6(a), we observe a notable transition from small to medium sizes, where averaged performance in three different datasets increases by 439.06%.
+The shift is more apparent under the logit difference metric (See Section 3.8 and Figure 6(b)).
+Given these results, we select GPT-2 Medium for circuit discovery and subsequently test the compatibility with different model sizes and architectures (Section 3.7 and 3.8).
+3.1 Transitive Reasoning Mechanisms
 Middle-term corruption reveals information flow relevant to the transitive property.
@@
 To verify this hypothesis, we further investigate the attention weights between $[p]$ and $[m_{2}]$.
-As expected, $[m_{2}]$ is the position most highly attended by $[p]$, with an average attention weight of $0.15\pm 0.07$ (Figure 4(e)).
+As expected, $[m_{2}]$ is the position most highly attended by $[p]$, with an average attention weights of $0.15\pm 0.07$ (Figure 3(e)).
 These results suggest that information is indeed moved from $[m_{2}]$ to $[p]$ on the residual stream subspace, with $\mathbf{h}_{11.10}$ playing a crucial role in this mechanism.
@@
 We notice that at position $[p]$, the model can observe the complete AAA-1 syllogistic structure (Figure 2(a)) with middle-term duplication.
-We hypothesize that this structural information for reasoning is collected at one position, given that information refinement occurs at the specific position such as last token (Hanna, Liu, and Variengien 2024; Stolfo, Belinkov, and Sachan 2023).
-To verify this, we trace the information flow of head $\mathbf{h}_{11.10}$ using path patching (Wang et al.
-2023).
-Our results show that $\mathbf{h}_{11.10}$ operates based on several induction heads (Elhage et al.
-2021) ($\mathbf{h}_{5.8}$, $\mathbf{h}_{6.1}$, $\mathbf{h}_{6.15}$ and $\mathbf{h}_{7.2}$) formed at $[m_{2}]$.
+We hypothesize that this structural information for reasoning is collected at one position, given that information refinement occurs at the specific position such as last token (Hanna et al., 2024; Stolfo et al., 2023).
+To verify this, we trace the information flow of head $\mathbf{h}_{11.10}$ using path patching (Wang et al., 2023).
+Our results show that $\mathbf{h}_{11.10}$ operates based on several induction heads (Elhage et al., 2021) ($\mathbf{h}_{5.8}$, $\mathbf{h}_{6.1}$, $\mathbf{h}_{6.15}$ and $\mathbf{h}_{7.2}$) formed at $[m_{2}]$.
 These heads attend to the $[[m_{1}]+1]$ token due to the $m_{1}\equiv m_{2}$ conditioned matching operation, likely containing $m_{1}$-related information from previous token heads.
 We conclude, therefore, that $m_{1}$ and $m_{2}$ information are aggregated at position $[p]$.
-Additional details of the path patching results are available in the Appendix F.
+Additional details of the path patching results are available in the Appendix G.
 Suppressive mechanism is revealed through Logit Lens.
 To better understand the internal mechanism occurring at the $[p]$ position, we investigate attention head $\mathbf{h}_{11.10}$.
-Having previously examined the attention pattern (composed of attention weights), we now focus on the attention value and output by analyzing the OV circuit ($W_{V}W_{O}$(Elhage et al.
-2021) using the logit lens method (Nostalgebraist 2020).
+Having previously examined the attention pattern (composed of attention weights), we now focus on the attention value and output by analyzing the OV circuit ($W_{V}W_{O}$(Elhage et al., 2021) using the logit lens method (Nostalgebraist, 2020).
 Interestingly, the result in Figure 3(c) shows a clear negative diagonal pattern suggesting that $\mathbf{h}_{11.10}$ strongly suppresses the logit when attending to the same token as the corresponding output.
@@
 For simplicity, we name head $\mathbf{h}_{11.10}$ as $m$-suppression head.
-Localization of Term-Related Information Flow
+3.2 Term-Related Information Flow
 Key information is moved from term-specific positions to the last position.
@@
 In the residual stream patching results (Figure 3(d)), we observe the highest positive score at the $[p]$ position, while negative scores are most prominent at the $[m_{1}]$ and $[m_{2}]$ positions, indicating that the information residing in these positions indeed contributes to their corresponding token prediction.
-This observation aligns with the importance of token embedding information at their respective positions given iterative refinement on the residual stream (Simoulin and Crabbé 2021).
+This observation aligns with the importance of token embedding information at their respective positions given iterative refinement on the residual stream (Simoulin and Crabbé, 2021).
 A closer examination of the last token position reveals that the negative effect propagates from relatively early layers (approximately layer 10 onwards), yet positive effects are from later layers (approximately layer 14 onwards), incurring a positive shift of the model’s prediction from $m$ to $p$.
@@
 Given the findings that the information for prediction resides in term-specific positions, we trace which attention heads transfer the information from each term position ($[p]$, $[m_{1}]$ and $[m_{2}]$) to the last token.
-We call these heads mover heads as the existence of “positive or negative mover heads” that carry or suppress information from the specific token position to the last token position (Wang et al.
-2023; García-Carrasco, Maté, and Carlos Trujillo 2024).
-To discover the source of information flow, we investigate the heads with high attention value weight ($W_{V}$) patching score and identify 9 notable mover heads (see Appendix G for details of mover heads localization).
+We call these heads mover heads as the existence of “positive or negative mover heads” that carry or suppress information from the specific token position to the last token position (Wang et al., 2023; García-Carrasco et al., 2024).
+To discover the source of information flow, we investigate the heads with high attention value-based patching score and identify 9 notable mover heads (see Appendix H for details of mover heads localization).
 By further investigation of attention patterns (Figure 3(f)–(h)), $\mathbf{h}_{14.14},\mathbf{h}_{15.14}$ and $\mathbf{h}_{18.12}$ are classified as positive copy heads attending highly to $[p]$, and $\mathbf{h}_{19.1}$ is a positive suppression head by attending to $[m_{2}]$ and $[s]$.
@@
 […]
-Summary
-Overall, we found that the internal mechanisms to arrive at the correct conclusion involve the following phases:
-  1. 1.
-    Long Induction: Early layers exhibit biases towards the wrong conclusion due to long-range repetition of the first premise “All A are B”.
-  2. 2.
-    Duplication: Induction heads aggregate information about duplicated middle terms in the premises.
-  3. 3.
-    Suppression: The model aggregates and inhibits middle-term information (i.e.,“B”) suppressing the long induction mechanism.
-  4. 4.
-    Mover: Token-specific information is propagated to the last token position. The process culminates in the prediction shift from B’ to the correct token ‘C’.
-Therefore, the discovered circuit is mainly characterized by an internal error correction mechanism.
-Interestingly, this mechanism is different from the way humans would reason on syllogistic arguments through the use of abstract logical primitives and inference rules.
-Circuit Evaluation
+3.3 Summary
+Overall, we found that the mechanisms to derive correct conclusion involve the following phases:
+(1) Long Induction:
+Early layers exhibit biases towards the wrong conclusion due to long-range repetition of the first premise “All A are B”.
+(2) Duplication:
+Induction heads aggregate information about duplicated middle terms in the premises.
+(3) Suppression:
+The model aggregates and inhibits middle-term information (i.e.,‘B’) suppressing the long induction mechanism.
+(4) Mover:
+Token-specific information is propagated to the last token position. The process culminates in the prediction shift from ‘B’ to the correct token ‘C’.
+Therefore, these results suggest that the discovered circuit is mainly characterized by an internal error correction mechanism.
+3.4 Circuit Evaluation
 To evaluate the comprehensive correctness of the circuit, we assess necessity and sufficiency via the ablation method described in section 2.
@@
 The ablation result demonstrates that both types of perturbations do not undermine the sufficiency and necessity property of the circuit.
-Circuit Transferability
+3.5 Circuit Transferability
 The circuit is necessary for non-symbolic syllogistic arguments, yet not sufficient for belief-inconsistent ones.
-We present ablation results for the two generated non-symbolic datasets: belief-consistent (Figure 6(a)) and belief-inconsistent (Figure 6(b)) (See Appendix H for the intervention results).
+We present ablation results for the two generated non-symbolic datasets: belief-consistent (Figure 5(a)) and belief-inconsistent (Figure 5(b)) (See Appendix I for the intervention results).
 Notably, the symbolic circuit proves necessary for both types of non-symbolic inputs, suggesting that the logic derived from symbolic syllogisms remains essential even when natural words are substituted.
@@
 This drastic difference supports our hypothesis that the knowledge acquired during pre-training corrupts the content-independent reasoning circuit identified on the symbolic set with additional attention heads.
-Generalization to Different Syllogistic Schemes
+3.6 Generalization to Syllogistic Schemes
 Mood-Figure C1 C2 C3 Acc
@@
 We further aim to understand whether the symbolic circuit is specific to the AAA-1 syllogism (Barbara).
-To this end, we extend our experiment to encompass all 15 unconditionally valid syllogisms (see Appendix A and I for details about the syllogistic schemes and more experimental results).
-To evaluate the transferability to all 15 schemes, we verify three main conditions, reported in Table 2: (C1) Necessity, (2) Sufficiency, and (C3) Positive Logit Difference.
+To this end, we extend our experiment to encompass all 15 unconditionally valid syllogisms (see Appendix A and J for details about the syllogistic schemes and more experimental results).
+To evaluate the transferability to all 15 schemes, we verify three main conditions, reported in Table 2: (C1) Necessity, (C2) Sufficiency, and (C3) Positive Logit Difference.
 We also measure the accuracy of the completion task for each syllogism.
@@
 Notably, all three conditions are satisfied for the affirmative syllogisms (AII-3, IAI-3, IAI-4), for which the model achieves an accuracy above 0.6.
-Generalization to Different Model Sizes
-Finally, we expand our analysis to different sizes of GPT-2 (small, large and XL), assuming core mechanisms would transfer across models trained on the same dataset.
+3.7 Generalization to Different Sizes
+Next, we expand our analysis to different sizes of GPT-2 (small, large and XL).
 Here, the intervention results show similar suppression mechanism patterns and information flow across all models.
-However, the residual stream of GPT-2 small in the all-term corruption setup is reversed due to its low downstream accuracy (see Appendix J).
+However, the residual stream of GPT-2 small in the all-term corruption setup is reversed due to its low downstream accuracy (see Appendix K for additional details).
 Moreover, as the model size increases, we found evidence that the contribution of attention heads becomes more complex.
-We hypothesize this might be caused by a stronger impact of world knowledge with increasing size, as suggested by the decrease in accuracy on the symbolic dataset and the increase on the non-symbolic set observed for the XL model (See Figure 3).
-Related Work
-Mechanistic circuit analysis has emerged as a promising approach to interpreting the internal mechanisms of Transformers (Olah et al.
-2020; Nanda et al.
-2023; Olsson et al.
-2022; Wang et al.
-2023; García-Carrasco, Maté, and Carlos Trujillo 2024; Hanna, Liu, and Variengien 2024).
-Existing approaches investigating internal reasoning mechanisms mainly focus on math-related tasks, elucidating the information flow for answering mathematical questions (Stolfo, Belinkov, and Sachan 2023) and examining arithmetic operations (Quirke and Barez 2024).
+We hypothesize this might be caused by a stronger impact of world knowledge with increasing size, as suggested by the decrease in accuracy on the symbolic dataset and the increase on the non-symbolic set observed for the XL model (See Figure 6(a)–(b)).
+3.8 Generalization to Different Models
+We further extend our investigation to assess whether a similar reasoning mechanism can be identified across different LMs.
+Specifically, we evaluate the circuit on: Pythia with 70M, 160M, 410M, and 1B parameters (Biderman et al., 2023); Qwen 2.5 with 0.5B and 1.5B parameters (Yang et al., 2024); Llama 3.2 with 1B parameters (Grattafiori et al., 2024); and additionally, GPT-2-medium after being fine-tuned on argumentative texts for critical thinking (Betz et al., 2021).
+Our results reveal distinct grouping patterns.
+In relatively small models—such as Pythia 70M and 160M—we observe erratic activation patterns that mirror the limitations seen in under-scaled models like GPT-2 small.
+In contrast, models with more than 0.5B parameters (namely, Pythia 410M, Pythia 1B, and Llama 3.2 1B) exhibit a compatible pattern in terms of information flow and suppressive mechanisms.
+Qwen exhibits a variant suppressive mechanism, with the suppression head operating at the final token position rather than $[p]$.
+Meanwhile, the fine-tuned GPT-2 medium model yields activation patterns that are nearly indistinguishable from its pre-trained counterpart, suggesting that the core reasoning circuit originates from pre-training rather than fine-tuning.
+Overall, our results indicate that while the core reasoning circuit generalizes effectively across sufficiently scaled models, subtle architectural nuances can lead to distinct but comparable internal dynamics (See Appendix L for the result details).
+4 Related Work
+Mechanistic circuit analysis has emerged as a promising approach to interpreting the internal mechanisms of Transformers (Olah et al., 2020; Nanda et al., 2023; Olsson et al., 2022; Wang et al., 2023; García-Carrasco et al., 2024; Hanna et al., 2024).
+Existing approaches investigating internal reasoning mechanisms mainly focus on math-related tasks, elucidating the information flow for answering mathematical questions (Stolfo et al., 2023) and examining arithmetic operations (Quirke and Barez, 2024).
 Most pertinent to our work, Wiegreffe et al.
 (2024) provides a mechanistic interpretation of multiple-choice question answering, investigating attention head-level patterns.
-In general, our mechanistic analysis complements recent work investigating the challenges in processing reasoning arguments that contradict established beliefs and whether the reasoning in LMs stems from internalized rules or memorized content (Ando et al.
-2023; Yu, Merullo, and Pavlick 2023; Wu et al.
-2024; Talmor et al.
-2020; Wu et al.
-2024; Kassner, Krojer, and Schütze 2020; Haviv et al.
-2023; Feldman 2020; Monea et al.
-2024; Yu, Merullo, and Pavlick 2023; Wallat, Singh, and Anand 2020).
-To address this question, different mechanistic approaches have been adopted to localizing factual associations (Meng et al.
-2022; Geva et al.
-2023; Dai et al.
-2022) or assessing conditions for generalisation (Wang et al.
-2024).
-However, to the best of our knowledge, we are the first to investigate content-independent mechanisms for syllogistic reasoning.
-Conclusion
+In general, our mechanistic analysis complements recent work investigating the challenges in processing reasoning arguments that contradict established beliefs and whether the reasoning in LMs stems from internalized rules or memorized content (Ando et al., 2023; Yu et al., 2023; Wu et al., 2024; Talmor et al., 2020; Wu et al., 2024; Kassner et al., 2020; Haviv et al., 2023; Feldman, 2020; Monea et al., 2024; Yu et al., 2023; Wallat et al., 2020; Eisape et al., 2024).
+To address this question, different mechanistic approaches have been adopted to localizing factual associations (Meng et al., 2022; Geva et al., 2023; Dai et al., 2022) or assessing conditions for generalization (Wang et al., 2024).
+However, to the best of our knowledge, we are the first to investigate mechanisms for syllogistic reasoning.
+5 Conclusion
 In this study, we have provided a comprehensive mechanistic interpretation of syllogistic reasoning within auto-regressive language models.
 Overall, our findings suggest that LMs indeed learn transferable content-independent reasoning mechanisms but that, at the same time, such mechanisms might be contaminated and suppressed by the same specific world knowledge acquired during pre-training.
-It is important to acknowledge some of the limitations in our study: (1) our analysis is conducted predominantly on the transitive property and term-specific information and does not consider the full spectrum of the reasoning dynamics that might appear in more complex scenarios.
-(2) Our experimentation is confined to a specific reasoning template that lacks the complexity and noise of real-world problems due to the intricate level of granularity and the computational complexity required in circuit analysis.
-Despite these limitations, we believe our findings could offer valuable insights into the reasoning mechanisms adopted by auto-regressive language models and lay a solid foundation for future research in this field.
+6 Limitations
+It is important to acknowledge some of the limitations of our study.
+First, our analysis is conducted predominantly on the transitive property and term-specific information and does not consider the full spectrum of the reasoning dynamics that might appear in more complex scenarios which are hard to model via causal intervention techniques.
+Second, Our experimentation is confined to specific syllogistic reasoning templates due to the intricate level of granularity and the computational complexity required in circuit analysis.
+Despite these limitations, however, we believe our findings could offer valuable insights into the reasoning mechanisms adopted by auto-regressive language models for reasoning along with existing biases and lay a solid foundation for future research in this field.
 Acknowledgements
-This work was partially funded by the Swiss National Science Foundation (SNSF) project NeuMath: Neural Discourse Inference over Mathematical Texts (200021_204617).
-Appendix
-This document serves as a technical appendix to the paper titled A Mechanistic Interpretation of Syllogistic Reasoning in Auto-Regressive Language Models.
-It provides additional theoretical background, experimental details, and supplementary results that support the main findings presented in the original paper.
-Appendix A A. Unconditionally Valid Syllogism Schemes
+This work was funded by the Swiss National Science Foundation (SNSF) project “NeuMath” (200021_204617), by the CRUK National Biomarker Centre, and supported by the Manchester Experimental Cancer Medicine Centre and the NIHR Manchester Biomedical Research Centre.
+Appendix A Unconditionally Valid Syllogism Schemes
 We list all 15 unconditionally valid syllogisms (Table 3).
@@
 Some B are A. No C are B. Some A are not C.
-Table 3: 15 Unconditionally valid syllogisms schemes.
+Table 3: 15 Unconditionally valid syllogism schemes.
 The table lists the syllogisms by their traditional names, moods, and figures, with formalised logical expressions on the first line and corresponding natural language examples on the second line.
-The minor premise (m) is presented before the major premise (M) to emphasise the transitive property inherent in these forms.
-Appendix B B. Dataset Generation
-B.1. Symbolic
+The minor premise (m) is presented before the major premise (M) to emphasise the transitive property in AAA-1 syllogism.
+Appendix B Dataset Generation
+B.1 Symbolic
 The symbolic dataset comprises sentences where all terms in the premises are represented by abstract symbols (uppercase alphabet letters).
 From the set of all 26 uppercase alphabets, three-letter triples (e.g., A, B and C) are randomly sampled.
-For each triple, six permutated prompts and label pairs are generated following a template designed to minimise latent semantic interference among alphabet symbol tokens.
-B.2. Non-symbolic
-The non-symbolic datasets are constructed based on GenericsKB (Bhakthavatsalam, Anastasiades, and Clark 2020), a resource that provides a foundation for evaluation of sentence veracity with associated truthfulness scores (0-1).
+For each triple, six permutated prompts and label pairs are generated following syllogism templates designed to minimise latent semantic interference among alphabet symbol tokens.
+B.2 Non-symbolic
+The non-symbolic datasets are constructed based on GenericsKB (Bhakthavatsalam et al., 2020), a resource that provides a foundation for the evaluation of sentence veracity with associated truthfulness scores (0-1).
 The dataset construction process involves the following steps:
-  • Extraction: Select generic sentences with a truthfulness score of 1 based on GenericsKB (Bhakthavatsalam, Anastasiades, and Clark 2020), specifically those in the form A are B, using regular expression.
+  • Extraction: Select generic sentences with a truthfulness score of 1 based on GenericsKB (Bhakthavatsalam et al., 2020), specifically those in the form A are B, using regular expression.
   • Syllogism Construction: We form universal affirmative syllogism arguments (Barbara) based on a template by chaining sentences where the predicate of one sentence logically matches the subject of another.
   • Constraints: We exclude terms tokenised into multiple tokens to maintain consistency in comparison with the symbolic dataset, which is essential for coherent circuit analysis.
-  • Classification: To address issues of partial inclusion and syntactic ambiguity in constructed syllogism arguments, we employ GPT-4 (OpenAI et al.
-2024) to classify whether arguments contain only truthful premises and whether the middle-terms are syntactically and semantically equivalent.
+  • Classification: To address issues of partial inclusion and syntactic ambiguity in constructed syllogism arguments, we employ GPT-4 (OpenAI et al., 2024) to classify whether arguments contain only truthful premises and whether the middle-terms are syntactically and semantically equivalent.
 This step helps us classify a belief-consistent dataset and a belief-inconsistent dataset.
@@
 This process ensures that our non-symbolic dataset maintains logical equivalence with the symbolic dataset while incorporating meaningful semantic real-word concepts.
-B.3. Data Statistics
+B.3 Data Statistics
 We organise data statistics for generated datasets (Table 4).
@@
 Table 4: Summary of Dataset Statistics. SYM refers to the symbolic dataset, BC refers to the non-symbolic belief-consistent dataset, and BI refers to the non-symbolic belief-inconsistent dataset.
-Appendix C C. Embedding Space Analysis Details
-One well-known method for analysing hidden activation in transformer language models is projecting activation into embedding space, the so-called logit lens (Nostalgebraist 2020; Elhage et al.
-2021; Geva et al.
-2022).
-In particular, we focus on an input-agnostic approach (Dar et al.
-2023; Hanna, Liu, and Variengien 2024), utilising both unembedding ($\mathbf{W}_{U}$) and embedding ($\mathbf{W}_{E}$) matrices to construct a $\mathbb{R}^{|V|\times|V|}$ matrix, where $|V|$ represents the model’s vocabulary size.
-In this study, we employ the OV circuit (Elhage et al.
-2021) of attention heads, formed by attention value and output weights ($W_{V}W_{O}$), to understand the general behaviour of how source tokens influence output logits.
-The OV circuit-based logit lens for attention head $\mathbf{h}$ is formalised as $\mathbf{W}_{E}\mathbf{W}^{\mathbf{h}}_{V}\mathbf{W}_{O}^{\mathbf{h}}\mathbf{W}_{U}\in\mathbb{R}^{|V|\times|V|}$.
-For simplicity and in line with established conventions (Elhage et al.
-2021; Dar et al.
-2023), this formulation ignores layer normalisation.
-Appendix D D. Circuit Ablation Method Details
+Appendix C Embedding Space Analysis Details
+One established method for analysing hidden activation in transformer language models is the logit lens, which projects activation into embedding space (Nostalgebraist, 2020; Elhage et al., 2021; Geva et al., 2022).
+We focus on an input-agnostic approach (Dar et al., 2023; Hanna et al., 2024), utilising both unembedding ($W_{U}$) and embedding ($W_{E}$) matrices to construct a $\mathbb{R}^{|V|\times|V|}$ matrix, where $|V|$ denotes the model’s vocabulary size.
+This study employs the OV circuit (Elhage et al., 2021) of attention heads, formed by attention value and output weights ($W_{V}W_{O}$), to understand how source tokens generally influence output logits.
+The OV circuit-based logit lens for attention head $\mathbf{h}$ is formalised as $W_{E}W^{\mathbf{h}}_{V}W_{O}^{\mathbf{h}}W_{U}\in\mathbb{R}^{|V|\times|V|}$.
+Following the previous works (Elhage et al., 2021; Dar et al., 2023), this formulation omits layer normalisation.
+Appendix D Circuit Ablation Method Details
 To measure the necessity and sufficiency of the circuit $\mathcal{C}$ in the model $\mathcal{M}$, we knock out attention heads in $\mathcal{H}$ from the model $\mathcal{M}$ and measure the average logit difference $\delta(p,m)$ along the batch.
@@
 $\mathbb{E}_{x\sim X}[\delta_{j}(p,m,\mathcal{C}_{j}^{*})]\quad\text{where}\quad\mathcal{C}_{j}^{*}=\mathcal{M}\setminus(\mathcal{H}\setminus H_{j})$
-Appendix E E. GPT-2 Model Architecture
+Appendix E Phase Transition
+In Figure 6(a–b), we observe a notable phase transition from small to medium models in accuracy and logit difference.
+Figure 6: Accuracy and logit difference on the syllogism completion task across all different sizes of GPT-2 for three different datasets.
+Appendix F GPT-2 Model Architecture
 We provide a self-contained concise overview of the GPT-2 model architecture, highlighting the main components and their mathematical relationships.
@@
 We refer the conventions of notation in Elhage et al.
-(2021) and (Geva et al.
-2023).
-Notation and Preliminaries
+(2021) and Geva et al.
+(2023).
+Notation
   • $x_{i}$ One-hot encoded vector representing the $i$-th token in the input sequence.
+  • $p_{i}$ One-hot encoded vector representing the $i$-th positional information.
+  • $r_{i}^{l}$ Residual stream vector representing the $i$-th token at layer $l$ in the input sequence.
   • $W_{E}^{\text{token}},W_{E}^{\text{pos}}$ Token and positional embedding matrices.
-  • $W_{Q},W_{K},W_{V}$ Query, key, and value weight matrices of attention heads.
+  • $A_{i}^{l}$ Output from the Multi-Head Self-Attention sublayer
+  • $M_{i}^{l}$ Output from Multi-Layer Perceptron (MLP) sublayer.
+  • $W_{Q},W_{K},W_{V},W_{O}$ Query, key, value, output weight matrices of attention heads.
   • $W^{\text{in}},W^{\text{out}}$ Input and output weight matrices of MLP feed-forward networks.
@@
 Each token in the input sequence is transformed into an embedded representation by combining token-specific and positional embeddings:
-$r_{i}^{0}=x_{i}W_{E}^{\text{token}}+x_{i}W_{E}^{\text{pos}}$
+$r_{i}^{0}=x_{i}W_{E}^{\text{token}}+p_{i}W_{E}^{\text{pos}}$
 This operation initializes the input for the transformer layers, where $r_{i}^{0}$ represents the initial embedded state of the $i$-th token.
@@
 $r_{i}^{l}=r_{i}^{l-1}+A_{i}^{l}+M_{i}^{l}$
-where $r_{i}^{l}$ represents the state of the $i$-th token at layer $l$, and $A_{i}^{l}$, $M_{i}^{l}$ are outputs from the Multi-Head Self-Attention (MHSA) and Multi-Layer Perceptron (MLP) sublayers.
-Multi-Head Self-Attention (MHSA).
+Multi-Head Self-Attention Sublayer.
 This sublayer processes information by applying self-attention mechanisms across multiple ‘heads’ of attention, enabling the model to capture various aspects of the input data, where $\sigma$ represent the non-linearity function:
-$A_{i}^{l}=\sum_{h\in H}\left(\sigma\left(\frac{(r_{i}^{l,h}W^{h}_{Q})(r_{i}^{l,h}W^{h}_{K})^{T}}{\sqrt{d_{K}}}\right)(r_{i}^{l,h}W^{h}_{V})\right)W^{h}_{O}$
+$A_{i}^{l}=\sum_{h\in H}\left(\sigma\left(\frac{(r_{i}^{l}W^{h}_{Q})(r_{i}^{l}W^{h}_{K})^{T}}{\sqrt{d_{K}}}\right)(r_{i}^{l}W^{h}_{V})\right)W^{h}_{O}$
 Each head computes a weighted sum of all tokens’ transformed states, focusing on different subsets of sequence information.
-Multi-Layer Perceptron (MLP).
+Multi-Layer Perceptron Sublayer.
 Each token’s representation is further processed in a position-wise manner by the MLP sublayer:
@@
 The MLP modifies each token’s state locally, enhancing its ability to process information.
-Layer Normalisation.
-Before processing by MHSA and MLP, each token’s state is normalised to stabilise learning and reduce training time:
+Layer Normalization.
+Before processing by self-attention and MLP processing, each token’s state is normalized to stabilise learning and reduce training time:
 $\text{LN}(r_{i}^{l})=\gamma\odot\frac{r_{i}^{l}-\mu_{i}^{l}}{\sqrt{(\sigma_{i}^{l})^{2}+\epsilon}}+\beta$
@@
 The predicted token is chosen based on the sampling method (e.g., greedy decoding) at the last position $N$.
-Appendix F F. Path Patching Details
-Path patching, an generalised version of activation patching, aims to compute direct effects among model components, rather than indirect effects diffused from intervened components (Wang et al.
-2023).
-This method involves freezing non-targeted activations during an initial forward pass, storing the corrupted targeted activation state, and then executing a subsequent forward pass with the targeted activation state substituted by the corrupted one.
+Appendix G Path Patching Details
+Path patching, an generalized version of activation patching, aims to compute direct effects among model components rather than indirect effects diffused from intervened components (Wang et al., 2023).
+This method involves freezing non-targeted activations during an initial forward pass, storing the targeted activation state in intervention, and then executing a subsequent forward pass with the targeted activation state substituted by the stored one.
 This approach isolates targeted component-to-component effects, eliminating non-relevant interactions.
@@
 Our findings, illustrated in Figure 7, reveal that $\mathbf{h}_{11.10}$ strongly operates based on several heads: $\mathbf{h}_{5.8}$, $\mathbf{h}_{6.1}$, $\mathbf{h}_{6.15}$, and $\mathbf{h}_{7.2}$.
-Manual investigation of attention patterns subsequently confirmed these as induction heads (Elhage et al.
-2021).
+Manual investigation of attention patterns subsequently confirmed these as induction heads (Elhage et al., 2021).
 Figure 7: Path patching results for (Top) symbolic inputs and non-symbolic inputs (Bottom) in the middle-term corruption setup.
 Note that negative heads are positively influencing heads, as we replace corrupted activation with clean ones in our noising approach.
-Appendix G G. Localising Mover Heads
-To efficiently classify the numerous attention heads (384 in GPT-2 medium) within the model, we calculate the Positional Patching Difference (PPD) to construct four-quadrant analysis.
+Appendix H Localizing Mover Heads
+To efficiently classify the numerous attention heads (384 in GPT-2 medium) within the model, we calculate the Positional Patching Difference (PPD) to construct a four-quadrant analysis:
+$\text{PPD}=|S[p]|-|S[m_{1}]+S[m_{2}]|$
 This method involves intervening in the attention value weights $W_{V}$ and constructs a distribution layout of heads, providing a systematic classification for quadrant groups.
-The layout is composed of the PPD score ($|S[p]|-|S[m_{1}]+S[m_{2}]|$) on the x-axis and the patching score for all sequence positions ($\mathcal{S}$) on the y-axis.
-Based on this PPD score quadrant layout analysis, we categorise heads according to their quadrant positions, revealing insights into their functional roles:
+The layout is composed of the PPD score on the x-axis and the patching score for all sequence positions ($\mathcal{S}$) on the y-axis.
+Based on this PPD score quadrant layout analysis, we categorize heads according to their quadrant positions, revealing insights into their functional roles:
   • (1)
@@
     Fourth quadrant (Negative Suppression Candidates): Heads with $\mathcal{S}<0$ but positive PPD, predominantly suppressing $[p]$-based information.
-The intervention of attention head values ($W_{V}$) is crucial because all-term corruption maintains the same positional alignment for all samples $(s,m_{1},m_{2},p)\rightarrow(s^{\prime},m_{1}^{\prime},m_{2}^{\prime},p^{\prime})$, resulting in minimal impact from attention head pattern interventions composed by attention query and key.
-Consequently, we can infer that the attention head output patching results primarily derive from the attention value activation.
-This property enables us to localise the source token positions from which the term information is moved by attention value patching.
-Formally, we denote the attention query, key, value, and output matrix activations as $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, $\mathbf{W}_{V}$, and $\mathbf{W}_{O}$ respectively.
-One attention head output can be expressed as $\sigma\left((\mathbf{R}\mathbf{W}_{Q})(\mathbf{R}\mathbf{W}_{K})^{T}\right)(\mathbf{R}\mathbf{W}_{V})\mathbf{W}_{O}$, where $\sigma$ represents the non-linearity function (softmax in GPT-2) applied in the attention pattern and $\mathbf{R}$ denotes the residual stream.
+The intervention of attention head values is selected because all-term corruption maintains the same positional alignment for all samples $(s,m_{1},m_{2},p)\rightarrow(s^{\prime},m_{1}^{\prime},m_{2}^{\prime},p^{\prime})$, resulting in minimal impact from attention head pattern interventions composed by attention query and key.
+Consequently, we can infer that the attention head output patching results primarily are derived from the attention value activation.
+This property enables us to localize the source token positions from which the term information is moved by attention value patching.
+Formally, when we denote the attention query, key, value, and output weight matrices as $W_{Q},W_{K},W_{V}$, and $W_{O}$ respectively, and the residual stream vector as $r$, one attention head output can be expressed as: $\sigma\left((rW_{Q})(rW_{K})^{T}\right)(rW_{V})W_{O}$, where $\sigma$ represents the non-linearity function (softmax in GPT-2) applied in the attention pattern.
 To include only influential heads in the circuit, we extracted outlier heads with an absolute patching score exceeding the threshold ($\tau$), defined by the mean and standard deviation of all heads’ scores as:
@@
 Our analysis identifies 9 heads for the symbolic circuit, as illustrated by red dots in Figure 8(a).
-$\mathbf{h}_{14.14}$, $\mathbf{h}_{15.14}$, and $\mathbf{h}_{18.12}$ are classified as positive copy heads, while $\mathbf{h}_{19.1}$ is grouped as a positive suppression head.
-$\mathbf{h}_{9.9}$, $\mathbf{h}_{11.1}$, $\mathbf{h}_{12.1}$, $\mathbf{h}_{17.2}$, and $\mathbf{h}_{23.10}$ are classified as negative copy heads.
-It is noteworthy that non-symbolic results reveal a more complex and noisy pattern in the mover localisation process (Figure 8(b)).
+$\mathbf{h}_{14.14}$, $\mathbf{h}_{15.14}$, and $\mathbf{h}_{18.12}$ are classified as positive copy candidates, while $\mathbf{h}_{19.1}$ is grouped as a positive suppression candidates.
+$\mathbf{h}_{9.9}$, $\mathbf{h}_{11.1}$, $\mathbf{h}_{12.1}$, $\mathbf{h}_{17.2}$, and $\mathbf{h}_{23.10}$ are classified as negative copy candidates.
+It is noteworthy that non-symbolic results reveal a more complex and noisy pattern in the mover localization process (Figure 8(b)).
 Figure 8: Distribution of attention heads in PPD-based quadrant analysis.
@@
 (b) Results for non-symbolic inputs.
-The x-axis represents normalised PPD scores, while the y-axis represents normalised attention value-based patching scores ($S$).
-Red dots indicate attention heads above the threshold ($\tau$) of patching score.
-Appendix H H. Non-symbolic Circuit Analysis
+The x-axis represents normalized PPD scores, while the y-axis represents normalized attention value-based patching scores ($S$).
+Red dots indicate attention heads above the threshold ($\tau$) of the patching score.
+Appendix I Non-symbolic Circuit Analysis
 We present the results of our non-symbolic circuit analysis (Figure 9).
@@
 For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.
-Appendix I I. Generalisability to Other Syllogisms
-To assess the generalisability of our model, we applied the circuit ablation method to all 15 unconditionally valid syllogisms (Table 3) and the result is shown in Figure 10.
+Appendix J Generalizability to Other Syllogisms
+To assess the generalizability of our model, we applied the circuit ablation method to all 15 unconditionally valid syllogisms (Table 3), and the result is shown in Figure 10.
 Unconditionally valid syllogisms are logical arguments that maintain their validity irrespective of the truth values of their premises.
 The validity of these syllogisms is determined solely by their logical form, which is characterised by mood and figure combinations.
-The mood of a syllogism is defined by the arrangement of four proposition types (”All (A)”, ”No (E)”, ”Some (I)”, ”Some … not (O)”) across its two premises and conclusion.
+The mood of a syllogism is defined by the arrangement of four proposition types ("All (A)", "No (E)", "Some (I)", "Some … not (O)") across its two premises and conclusion.
 Conversely, the figure of a categorical syllogism is determined by the structural arrangement of terms within the constituent sentences.
@@
 Figure 10: Necessity and sufficiency performance results from circuit ablation method for all unconditionally valid syllogistic forms. Labels (e.g., AII-3) denote mood and figure combinations.
-Appendix J J. Generalisability Across Model Sizes
-To assess the generalisability of our findings across different model capacities, we extended our symbolic circuit analysis to various GPT-2 model sizes.
+Model S BC BI
+Group 1: Erratic Pattern
+Pythia-70M 0.19 0.02 0.02
+Pythia-160M 0.00 0.12 0.21
+Group 2: Compatible Pattern
+Pythia-410M 0.99 0.40 0.51
+Pythia-1B 0.30 0.31 0.43
+Llama-3.2-1B 1.00 0.81 0.70
+Group 3: Variant Pattern
+Qwen2.5-0.5B 1.00 0.96 0.90
+Qwen2.5-1.5B 1.00 0.99 0.90
+Table 5: Accuracy comparison of different models across three evaluation categories: Symbolic (S), Belief-Consistent (BC), and Belief-Inconsistent (BI).
+Group 1 models exhibit erratic activation patterns, failing to establish a stable reasoning mechanism.
+Group 2 models demonstrate a compatible reasoning circuit with consistent information flow and suppression.
+Group 3, particularly the Qwen models, exhibit a functional but distinct suppression mechanism, where suppression is applied at the final token rather than the $[p]$ position.
+Appendix K Generalizability Across Model Sizes
+Our approach is guided by the universality hypothesis, which suggests that similar representational patterns and circuits emerge across models trained on the same dataset (Olah et al., 2020; Kornblith et al., 2019; Li et al., 2015).
+We assume that core mechanisms should persist across these different model sizes.
 Figure 11 presents the results of this comparative analysis.
@@
 For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.
+Appendix L Generalizability Across Different Models
+Figures 12 and 13 represent the results of applying the same circuit analysis in different model families.
+The $m$-suppression heads for each model were manually selected based on their likelihood, determined by patching scores and spatial positioning within the activation pattern.
+The corresponding model accuracies in the three datasets (Symbolic, Belief-Consistent, and Belief-Inconsistent) are presented in Table 5.
+Figure 12: Comprehensive results of the symbolic circuit analysis across different series of models (Pythia).
+(a) Attention output patching results and (b) residual stream patching results in the middle-term intervention setup.
+(c) OV circuit logit lens results for $m$-suppression head, with input and output comprising 26 uppercase letters.
+(d) Residual stream patching results in the all-term corruption setup.
+For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.
+Figure 13: Comprehensive results of the symbolic circuit analysis across different series of models (Qwen2.5, Llama3.2 and fine-tuned GPT-2 medium).
+(a) Attention output patching results and (b) residual stream patching results in the middle-term intervention setup.
+(c) OV circuit logit lens results for $m$-suppression head, with input and output comprising 26 uppercase letters.
+(d) Residual stream patching results in the all-term corruption setup.
+For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.
+Appendix M Reproducibility Statement
+Our study is conducted using TransformerLens (Nanda and Bloom, 2022) on an Nvidia A100 GPU with 80GB of memory.
+The dataset and code to reproduce our experiments are available online to encourage future work in the field [^1] http://anonynous_url.com.
+Appendix N Statistical Significance
+For all experiments, we use 90 samples each for both symbolic and non-symbolic arguments to balance analytical depth with computational efficiency.
+This sample size sufficiently yields statistically significant activation patching results ($p<0.05$).

### v3

@@
-A Mechanistic Interpretation of Syllogistic Reasoning in Auto-Regressive Language Models
+Reasoning Circuits in Language Models: A Mechanistic Interpretation of Syllogistic Inference
 Abstract
-Recent studies on logical reasoning in Language Models (LMs) have sparked a debate on whether they can learn systematic reasoning principles during pre-training or merely exploit superficial patterns in the training data.
-This paper presents a mechanistic interpretation of syllogistic reasoning in LMs to advance the understanding of internal dynamics.
-Specifically, we present a methodology for circuit discovery aimed at interpreting content-independent reasoning mechanisms.
+Recent studies on reasoning in language models (LMs) have sparked a debate on whether they can learn systematic inferential principles or merely exploit superficial patterns in the training data.
+To understand and uncover the mechanisms adopted for formal reasoning in LMs, this paper presents a mechanistic interpretation of syllogistic inference.
+Specifically, we present a methodology for circuit discovery aimed at interpreting content-independent and formal reasoning mechanisms.
 Through two distinct intervention methods, we uncover a sufficient and necessary circuit involving middle-term suppression that elucidates how LMs transfer information to derive valid conclusions from premises.
-Furthermore, we investigate how belief biases manifest in syllogistic reasoning, finding evidence of partial contamination from additional attention heads responsible for encoding commonsense and contextualized knowledge.
-Finally, we explore the generalization of the discovered mechanisms across various syllogistic schemes, model sizes and architectures, finding that the identified circuit is sufficient and necessary for the schemes on which the models achieve high downstream accuracy ($\geq$ 60%), and that the activation patterns apply to models of different families.
-Overall, our findings suggest that LMs indeed learn transferable content-independent reasoning mechanisms, but that, at the same time, such mechanisms do not involve generalizable and abstract logical primitives, being susceptible to contamination by the same world knowledge acquired during pre-training.
-A Mechanistic Interpretation of Syllogistic Reasoning in
-Auto-Regressive Language Models
-Geonhee Kim1, Marco Valentino3, André Freitas1,2,3 1Department of Computer Science, University of Manchester, UK 2National Biomarker Centre (NBC), CRUK Manchester Institute, UK 3 Idiap Research Institute, Switzerland
+Furthermore, we investigate how belief biases manifest in syllogistic inference, finding evidence of partial contamination from additional attention heads responsible for encoding commonsense and contextualized knowledge.
+Finally, we explore the generalization of the discovered mechanisms across various syllogistic schemes, model sizes and architectures.
+The identified circuit is sufficient and necessary for syllogistic schemes on which the models achieve high accuracy ($\geq$ 60%), with compatible activation patterns across models of different families.
+Overall, our findings suggest that LMs learn transferable content-independent reasoning mechanisms, but that, at the same time, such mechanisms do not involve generalizable and abstract logical primitives, being susceptible to contamination by the same world knowledge acquired during pre-training.
+Reasoning Circuits in Language Models:
+A Mechanistic Interpretation of Syllogistic Inference
+Geonhee Kim1, Marco Valentino2,3, André Freitas1,2,4 1Department of Computer Science, University of Manchester, UK 2Idiap Research Institute, Switzerland 3School of Computer Science, University of Sheffield, UK 4National Biomarker Centre, CRUK-MI, University of Manchester, UK kimgeonhee317@gmail.com, ac4mv@sheffield.ac.uk, andre.freitas@idiap.ch
 1 Introduction
@@
 The process culminates in the prediction shift from ‘B’ to the correct token, ‘C.’
-Transformer-based Language Models (LMs) have led to remarkable results across various natural language processing tasks (Radford et al., 2018, 2019; Brown et al., 2020; OpenAI et al., 2024; Jason et al., 2022; Bubeck et al., 2023).
+Language models (LMs) have led to remarkable results across various natural language processing tasks (Radford et al., 2018, 2019; Brown et al., 2020; OpenAI, 2024; Jason et al., 2022; Bubeck et al., 2023).
 This success has catalyzed research interest in systematically exploring the reasoning capabilities emerging during pre-training (Clark et al., 2020).
-Recent findings suggest that logical reasoning abilities may emerge in large-scale models (Rae et al., 2022; Kojima et al., 2022; Wei et al., 2022) or through transfer learning on specialized datasets (Betz et al., 2021).
+Recent findings suggest that logical and formal reasoning abilities may emerge in large-scale models (et al., 2022; Kojima et al., 2022; Wei et al., 2022) or through transfer learning on specialized datasets (Betz et al., 2021).
 However, ongoing debates question whether these models apply systematic inference rules or reuse superficial patterns learned during pre-training (Talmor et al., 2020; Kassner et al., 2020; Wu et al., 2024).
-This controversy underscores the need for a deeper understanding of the low-level logical inference mechanisms in LMs (Rozanova et al., 2024, 2023, 2022).
+This controversy underscores the need for a deeper understanding of the low-level logical inference mechanisms in LMs (Rozanova et al., 2024, 2023, 2022; Yanaka et al., 2020).
 To improve our understanding of the internal mechanisms, this paper focuses on mechanistic interpretability (Olah et al., 2020; Nanda et al., 2023), aiming to discover the core circuit responsible for syllogistic reasoning.
@@
 Second, AAA-1 is a form of unconditionally valid syllogism independent of the premises’ truth condition (Holyoak and Morrison, 2005).
-This characteristic offers a deterministic and scalable task design, other than allowing us to investigate the disentanglement between reasoning and knowledge representation Wysocka et al.
-(2024); Bertolazzi et al.
+This characteristic offers a deterministic and scalable task design, other than allowing us to investigate the disentanglement between reasoning and knowledge representation Bertolazzi et al.
+(2024); Wysocka et al.
 (2025); Lampinen et al.
+(2024); Valentino et al.
 (2025).
-Through mechanistic interpretability techniques such as Activation Patching (Meng et al., 2022) and embedding space analysis (i.e., Logit Lens) (Nostalgebraist, 2020; Geva et al., 2022; Dar et al., 2023), we investigate the following main research questions: RQ1: How is the content-independent syllogistic reasoning mechanism internalized in LMs during pre-training?; RQ2: Are content-independent mechanisms disentangled from specific world knowledge and belief biases?; RQ3: Does the core reasoning mechanism generalize across model sizes and syllogistic schemes?
+Through mechanistic interpretability techniques such as Activation Patching (Meng et al., 2022) and embedding space analysis (i.e., Logit Lens) (Nostalgebraist, 2020; Geva et al., 2022; Dar et al., 2023), we investigate the following main research questions: RQ1: How is the content-independent syllogistic reasoning mechanism internalized in LMs during pre-training?; RQ2: Are content-independent mechanisms disentangled from specific world knowledge and belief biases?; RQ3: Does the core reasoning mechanism generalize across syllogistic schemes, different model sizes, and architectures?
 To answer these questions, we present a methodology that consists of 3 main stages.
@@
 1.
-The circuit analysis reveals that LMs develop specific inference mechanisms during pre-training, finding evidence supporting a three-stage mechanism for syllogistic reasoning: (1) naive recitation of the first premise; (2) suppression of duplicated middle-term information; and (3) mediation towards the correct output through the interplay of attention heads.
+The circuit analysis reveals that LMs develop specific inference mechanisms during pre-training, finding evidence supporting a three-stage mechanism for syllogistic reasoning: (1) naive recitation of the first premise; (2) suppression of duplicated middle-term information; and (3) mediation towards the correct output through the interplay of mover attention heads (see Figure 1).
 2.
@@
 3.
-We found that the identified circuit is sufficient and necessary for all the syllogistic schemes in which the model achieves high downstream accuracy ($\geq$ 60%).
+We found that the identified circuit is sufficient and necessary for all the unconditionally valid syllogistic schemes in which the model achieves high downstream accuracy ($\geq$ 60%) (see Appendix A for the list of schemes).
 This result suggests that LMs learn reasoning mechanisms that are transferable across different schemes.
 4.
-The intervention results on models with different architectures and sizes (i.e., GPT-2 (Radford et al., 2019), Pythia (Biderman et al., 2023), Llama (Grattafiori et al., 2024), Qwen (Yang et al., 2024)) show similar suppression mechanism patterns and information flow.
+The intervention results on models with different architectures and sizes (i.e., GPT-2 (Radford et al., 2019), Pythia (Biderman et al., 2023), Llama (et al., 2024), Qwen (Yang et al., 2024)) show similar suppression mechanism patterns and information flow.
 However, we found evidence that the contribution of attention heads becomes more complex with model sizes, further supporting the hypothesis of increasing contamination from external world knowledge.
@@
 The uncovered symbolic circuit is evaluated to determine whether the core reasoning schema operates independently from world knowledge or belief bias.
+Our study is conducted using TransformerLens (Nanda and Bloom, 2022) on an Nvidia A100 GPU with 80GB of memory.
+The dataset and code to reproduce our experiments are available online to encourage future work in the field [^1] https://github.com/neuro-symbolic-ai/Mechanistic-Interpretation-Syllogism.
 2 Methodology
@@
 Our study primarily examines the activation of residual streams and multi-head self-attention at the sequence level to trace the term information flow.
+Activation patching includes three model runs (Clean, Corrupted and Intervened) alongside a quantification process to measure the effect of the interventions (see Figure 2(1)–(4)).
 Given a masked syllogistic input $x=[\mathcal{P}_{1};\mathcal{P}_{2};\mathcal{C}\setminus\{p\}]$ as $(s,m_{1},m_{2},p)$, which can be read as “All $s$ are $m_{1}$.
@@
 (1) Clean Run.
-For the target input $(s,m_{1},m_{2},p)$, we record the baseline logit difference, $\delta_{+}(p,m)$ from the forward pass output of the model (Figure 2(1)).
+For the target syllogistic input $(s,m_{1},m_{2},p)$, we record the baseline logit difference, $\delta_{+}(p,m)$ from the forward pass output of the model (Figure 2(1)).
 (2) Corrupted Run.
@@
 Additional technical details can be found in Appendix C.
-Figure 3: Results of the symbolic circuit analysis.
-(a) Attention output patching and (b) residual stream patching obtained via middle-term intervention.
-(c) OV circuit logit lens results for attention head $\mathbf{h}_{11.10}$, with input and output comprising 26 uppercase letters.
-(d) Residual stream patching obtained via all-term corruption.
-(e)-(h) Average attention weights across the batch: (e) weights for token position $[p]$, and (f)-(h) weights for the last token position.
-For clarity, the symbol ‘-’ on the y-axis denotes the averaged values for tokens appearing between terms.
-2.3 Interventions
+Figure 3: Visualization of the symbolic reasoning circuit identified in the model.
+(a) Attention output patching and (b) residual stream patching using middle-term interventions highlight the role of attention head $\mathbf{h}_{11.10}$ in aggregating information from the duplicated middle term to the predicate token position $[p]$.
+(c) Logit lens visualization of the OV (output vector) of $\mathbf{h}_{11.10}$ across 26 uppercase letters reveals its suppressive effect on the attended token at $[p]$.
+(d) Residual stream patching with all-term corruption indicates that key reasoning information is concentrated at term token positions.
+(e)–(h) Average attention weights across the batch: (e) attention from token position $[p]$, attending strongly to $[m_{2}]$; (f)–(h) attention from the last token position, identifying which heads (mover heads) propagate term information to the output.
+For clarity, a dash (–) on the axis indicates the averaged values for tokens appearing between terms.
+2.3 Causal Interventions
 The choice of tokens to corrupt is a critical aspect of the experimental design, and it is essential to establish an appropriate type of intervention that aligns with the hypothesis being tested.
@@
 We hypothesize that if heads carry the information preferring $p$ over $m$ for the valid prediction, the logit difference will increase; otherwise, it will decrease.
-2.4 Circuit Ablation.
+2.4 Circuit Ablation and Evaluation
 Once we discover a relevant circuit $\mathcal{C}$, we evaluate its necessity and sufficiency using mean ablation.
@@
 Additional details are included in Appendix D.
-3 Empirical Evaluation
+3 Mechanistic Interpretation
 Empirical Setup.
-We select the LM for the circuit analysis based on a trade-off between model size and performance.
-To this end, we measure the accuracy of GPT-2 (Radford et al., 2019) with different sizes (117M, 345M, 762M, and 1.5B) on the syllogism completion task.
-Specifically, we aim to identify potential phase transition points where significant performance changes emerge (Jason et al., 2022; Kaplan et al., 2020).
-As shown in Figure 6(a), we observe a notable transition from small to medium sizes, where averaged performance in three different datasets increases by 439.06%.
-The shift is more apparent under the logit difference metric (See Section 3.8 and Figure 6(b)).
-Given these results, we select GPT-2 Medium for circuit discovery and subsequently test the compatibility with different model sizes and architectures (Section 3.7 and 3.8).
+We select the LM for circuit analysis based on a trade-off between model size and performance.
+To this end, we measure the accuracy of GPT-2 (Radford et al., 2019) at various sizes (117M, 345M, 762M, and 1.5B) on the syllogism completion task, aiming to identify potential phase transition points where performance changes significantly (Jason et al., 2022; Kaplan et al., 2020).
+We observe a marked transition from small to medium sizes, where average performance across three datasets increases by 439.06% (see Figure 6(a), Appendix E).
+The shift is even more pronounced under the logit difference metric (see Section 4.4 and Figure 6(b), Appendix E).
+Based on these findings, we select GPT-2 Medium—whose architecture is summarized in Appendix F—for circuit discovery, and subsequently evaluate generalization across model sizes and architectures (Sections 4.3 and 4.4).
 3.1 Transitive Reasoning Mechanisms
@@
 We hypothesize that this structural information for reasoning is collected at one position, given that information refinement occurs at the specific position such as last token (Hanna et al., 2024; Stolfo et al., 2023).
-To verify this, we trace the information flow of head $\mathbf{h}_{11.10}$ using path patching (Wang et al., 2023).
+To verify this, we employ path patching (Wang et al., 2023), a more selective variant of activation patching, to trace the information flow of head $\mathbf{h}_{11.10}$.
 Our results show that $\mathbf{h}_{11.10}$ operates based on several induction heads (Elhage et al., 2021) ($\mathbf{h}_{5.8}$, $\mathbf{h}_{6.1}$, $\mathbf{h}_{6.15}$ and $\mathbf{h}_{7.2}$) formed at $[m_{2}]$.
@@
 Additional details of the path patching results are available in the Appendix G.
-Suppressive mechanism is revealed through Logit Lens.
+A suppressive mechanism is revealed through Logit Lens.
 To better understand the internal mechanism occurring at the $[p]$ position, we investigate attention head $\mathbf{h}_{11.10}$.
@@
 We call these heads mover heads as the existence of “positive or negative mover heads” that carry or suppress information from the specific token position to the last token position (Wang et al., 2023; García-Carrasco et al., 2024).
-To discover the source of information flow, we investigate the heads with high attention value-based patching score and identify 9 notable mover heads (see Appendix H for details of mover heads localization).
-By further investigation of attention patterns (Figure 3(f)–(h)), $\mathbf{h}_{14.14},\mathbf{h}_{15.14}$ and $\mathbf{h}_{18.12}$ are classified as positive copy heads attending highly to $[p]$, and $\mathbf{h}_{19.1}$ is a positive suppression head by attending to $[m_{2}]$ and $[s]$.
-$\mathbf{h}_{9.9}$, $\mathbf{h}_{11.1}$, $\mathbf{h}_{12.1}$, $\mathbf{h}_{17.2}$ and $\mathbf{h}_{23.10}$ are negative copy heads giving a strong attention to non-$[p]$.
+To identify the sources of information flow, we apply attention value-based patching and isolate nine notable mover heads (see Appendix H for localization details).
+Among them, $\mathbf{h}_{14.14}$, $\mathbf{h}_{15.14}$, and $\mathbf{h}_{18.12}$ act as positive copy heads, attending strongly to $[p]$, while $\mathbf{h}_{19.1}$ functions as a positive suppression head, exhibiting high attention to $[m_{2}]$ and $[s]$ (see Figure 3(f)).
+In contrast, $\mathbf{h}_{9.9}$, $\mathbf{h}_{11.1}$, $\mathbf{h}_{12.1}$, $\mathbf{h}_{17.2}$, and $\mathbf{h}_{23.10}$ serve as negative copy heads, focusing primarily on $[m_{1}]$ (Figure 3(g)) or showing more diffuse pattern (Figure 3(h)).
 A long induction mechanism in early layer movers causes biases towards the middle term.
 […]
-3.3 Summary
-Overall, we found that the mechanisms to derive correct conclusion involve the following phases:
+3.3 Circuit Discovery Summary
+Overall, we found that the mechanisms for syllogistic inference involve the following phases:
 (1) Long Induction:
@@
 (4) Mover:
-Token-specific information is propagated to the last token position. The process culminates in the prediction shift from ‘B’ to the correct token ‘C’.
-Therefore, these results suggest that the discovered circuit is mainly characterized by an internal error correction mechanism.
-3.4 Circuit Evaluation
+Token-specific information is propagated to the last position. The process ends in the prediction shift from ‘B’ to the correct token ‘C’.
+These results suggest that the circuit is characterized by an internal error correction mechanism.
+Interestingly, this mechanism is different from the way human experts would reason on syllogistic arguments through the systematic application of abstract logical primitives and inference rules.
+4 Circuit Evaluation
+The circuit is sufficient and necessary for symbolic arguments.
 To evaluate the comprehensive correctness of the circuit, we assess necessity and sufficiency via the ablation method described in section 2.
@@
 Solid lines show numeric perturbation-based performance, dotted lines indicate quantifier perturbation-based performance in (b), and the dashed line shows the baseline logit difference without knockouts.
-Figure 5: Circuit ablation results for evaluating transferability on non-symbolic datasets: the belief-consistent (a) and the belief-inconsistent (b).
+These results validate the necessity, sufficiency, and robustness of the identified symbolic reasoning circuit.
+Figure 5: Circuit ablation results for evaluating transferability on non-symbolic datasets: (a) the belief-consistent and (b) the belief-inconsistent.
 The dashed line represents the baseline logit difference without knockouts.
+In both cases, the symbolic circuit is necessary, as ablating it reduces performance.
+However, the circuit is only sufficient for belief-consistent data, as recovery fails in belief-inconsistent settings, suggesting interference from belief-bias.
+The circuit is robust to superficial variations.
 We further verify the robustness of the symbolic circuit to superficial and semantic-preserving perturbations.
@@
 The ablation result demonstrates that both types of perturbations do not undermine the sufficiency and necessity property of the circuit.
-3.5 Circuit Transferability
-The circuit is necessary for non-symbolic syllogistic arguments, yet not sufficient for belief-inconsistent ones.
-We present ablation results for the two generated non-symbolic datasets: belief-consistent (Figure 5(a)) and belief-inconsistent (Figure 5(b)) (See Appendix I for the intervention results).
+4.1 Circuit Transferability to Concrete Arguments
+The circuit is necessary for non-symbolic arguments, yet not sufficient for belief-inconsistent ones.
+We present ablation results for the two generated non-symbolic datasets: belief-consistent (Figure 5(a)) and belief-inconsistent (Figure 5(b)).
 Notably, the symbolic circuit proves necessary for both types of non-symbolic inputs, suggesting that the logic derived from symbolic syllogisms remains essential even when natural words are substituted.
-Regarding sufficiency, while belief-consistent data show significant performance recovery, we observe an inability to restore original performance on the belief-inconsistent set despite an increasing trend.
+Regarding sufficiency, while belief-consistent data show significant performance recovery, we observe an inability to restore performance on the belief-inconsistent set despite an increasing trend.
 These results indicate that belief biases encoded in different attention heads may play an important role.
@@
 ‘Accuracy (Acc)’ shows performance on the completion task for each syllogism.
+The results show that the model achieves high downstream accuracy particularly for affirmative syllogisms.
 We further aim to understand whether the symbolic circuit is specific to the AAA-1 syllogism (Barbara).
-To this end, we extend our experiment to encompass all 15 unconditionally valid syllogisms (see Appendix A and J for details about the syllogistic schemes and more experimental results).
+To this end, we extend our experiment to encompass all 15 unconditionally valid syllogisms (see Appendix A and I for details about the syllogistic schemes and more experimental results).
 To evaluate the transferability to all 15 schemes, we verify three main conditions, reported in Table 2: (C1) Necessity, (C2) Sufficiency, and (C3) Positive Logit Difference.
@@
 Notably, all three conditions are satisfied for the affirmative syllogisms (AII-3, IAI-3, IAI-4), for which the model achieves an accuracy above 0.6.
-3.7 Generalization to Different Sizes
+4.3 Generalization to Model Sizes
 Next, we expand our analysis to different sizes of GPT-2 (small, large and XL).
 Here, the intervention results show similar suppression mechanism patterns and information flow across all models.
-However, the residual stream of GPT-2 small in the all-term corruption setup is reversed due to its low downstream accuracy (see Appendix K for additional details).
+However, the residual stream of GPT-2 small in the all-term corruption setup is reversed due to its low downstream accuracy (see Appendix J for additional details).
 Moreover, as the model size increases, we found evidence that the contribution of attention heads becomes more complex.
-We hypothesize this might be caused by a stronger impact of world knowledge with increasing size, as suggested by the decrease in accuracy on the symbolic dataset and the increase on the non-symbolic set observed for the XL model (See Figure 6(a)–(b)).
-3.8 Generalization to Different Models
+We hypothesize this might be caused by a stronger impact of world knowledge with increasing size, as suggested by the decrease in accuracy on the symbolic dataset and the increase on the non-symbolic set observed for the XL model (see Figure 6(a)–(b), Appendix E).
+4.4 Generalization to Model Families
 We further extend our investigation to assess whether a similar reasoning mechanism can be identified across different LMs.
-Specifically, we evaluate the circuit on: Pythia with 70M, 160M, 410M, and 1B parameters (Biderman et al., 2023); Qwen 2.5 with 0.5B and 1.5B parameters (Yang et al., 2024); Llama 3.2 with 1B parameters (Grattafiori et al., 2024); and additionally, GPT-2-medium after being fine-tuned on argumentative texts for critical thinking (Betz et al., 2021).
+Specifically, we evaluate the circuit on: Pythia with 70M, 160M, 410M, and 1B parameters (Biderman et al., 2023); Qwen 2.5 with 0.5B and 1.5B parameters (Yang et al., 2024); Llama 3.2 with 1B parameters (et al., 2024); and additionally, GPT-2-medium after being fine-tuned on argumentative texts for critical thinking (Betz et al., 2021).
 Our results reveal distinct grouping patterns.
@@
 In contrast, models with more than 0.5B parameters (namely, Pythia 410M, Pythia 1B, and Llama 3.2 1B) exhibit a compatible pattern in terms of information flow and suppressive mechanisms.
-Qwen exhibits a variant suppressive mechanism, with the suppression head operating at the final token position rather than $[p]$.
+Qwen exhibits a variant suppressive mechanism, with the suppression head operating at the last token position rather than $[p]$.
 Meanwhile, the fine-tuned GPT-2 medium model yields activation patterns that are nearly indistinguishable from its pre-trained counterpart, suggesting that the core reasoning circuit originates from pre-training rather than fine-tuning.
-Overall, our results indicate that while the core reasoning circuit generalizes effectively across sufficiently scaled models, subtle architectural nuances can lead to distinct but comparable internal dynamics (See Appendix L for the result details).
+Table 3 summarizes the accuracy of each model across the Symbolic, Belief-Consistent, and Belief-Inconsistent datasets (see further circuit analysis details in Appendix K).
+Overall, our results indicate that while the reasoning circuit generalizes effectively across sufficiently scaled models, subtle architectural nuances can lead to distinct but comparable dynamics.
+Model S BC BI
+Group 1: Erratic Pattern
+Pythia-70M 0.19 0.02 0.02
+Pythia-160M 0.00 0.12 0.21
+Group 2: Compatible Pattern
+Pythia-410M 0.99 0.40 0.51
+Pythia-1B 0.30 0.31 0.43
+Llama-3.2-1B 1.00 0.81 0.70
+Group 3: Variant Pattern
+Qwen2.5-0.5B 1.00 0.96 0.90
+Qwen2.5-1.5B 1.00 0.99 0.90
+Table 3: Accuracy comparison across different models: symbolic (S), belief-consistent (BC), and inconsistent (BI).
+Group 1 models exhibit erratic patterns, failing to establish a stable reasoning mechanism.
+Group 2 models exhibit a compatible reasoning circuit with consistent information flow and suppression.
+Group 3 exhibits a functional but distinct suppression mechanism, where suppression is applied at the last token position.
 5 Related Work
@@
 6 Conclusion
-In this study, we have provided a comprehensive mechanistic interpretation of syllogistic reasoning within auto-regressive language models.
-Overall, our findings suggest that LMs indeed learn transferable content-independent reasoning mechanisms but that, at the same time, such mechanisms might be contaminated and suppressed by the same specific world knowledge acquired during pre-training.
+In this study, we presented a comprehensive mechanistic interpretation of syllogistic reasoning in language models.
+By combining activation patching, embedding space analysis, and circuit ablation, we uncovered a structured error-correction mechanism characterized by the suppression of duplicated middle terms and the propagation of relevant information via mover heads.
+Our findings show that this circuit is necessary and sufficient for symbolic syllogistic reasoning.
+Moreover, we demonstrated that, while this reasoning mechanism generalizes across syllogistic schemes, model architectures and sizes, it remains susceptible to contamination from belief biases when instantiated with natural language inputs.
+Overall, our findings suggest that LMs learn transferable reasoning mechanisms but that, at the same time, such mechanisms might be contaminated and suppressed by the same world knowledge acquired during pre-training.
+Further studies will be required to understand how the discovered symbolic circuit can be effectively disentangled from world knowledge to enable systematic generalization in LMs’ reasoning.
 7 Limitations
@@
 Second, Our experimentation is confined to specific syllogistic reasoning templates due to the intricate level of granularity and the computational complexity required in circuit analysis.
-Despite these limitations, however, we believe our findings could offer valuable insights into the reasoning mechanisms adopted by auto-regressive language models for reasoning along with existing biases and lay a solid foundation for future research in this field.
+Despite these limitations, however, we believe our findings could offer valuable insights into the reasoning mechanisms adopted by language models for formal reasoning and related biases, laying a solid foundation for future research in this field.
 Acknowledgements
@@
 Table 4: 15 Unconditionally valid syllogism schemes.
-The table lists the syllogisms by their traditional names, moods, and figures, with formalised logical expressions on the first line and corresponding natural language examples on the second line.
-The minor premise (m) is presented before the major premise (M) to emphasise the transitive property in AAA-1 syllogism.
+The table lists the syllogisms by their traditional names, moods, and figures, with formalized logical expressions on the first line and corresponding natural language examples on the second line.
+The minor premise (m) is presented before the major premise (M) to emphasize the transitive property in AAA-1 syllogism.
 Appendix B Dataset Generation
@@
   • Syllogism Construction: We form universal affirmative syllogism arguments (Barbara) based on a template by chaining sentences where the predicate of one sentence logically matches the subject of another.
-  • Constraints: We exclude terms tokenised into multiple tokens to maintain consistency in comparison with the symbolic dataset, which is essential for coherent circuit analysis.
-  • Classification: To address issues of partial inclusion and syntactic ambiguity in constructed syllogism arguments, we employ GPT-4 (OpenAI et al., 2024) to classify whether arguments contain only truthful premises and whether the middle-terms are syntactically and semantically equivalent.
+  • Constraints: We exclude terms tokenized into multiple tokens to maintain consistency in comparison with the symbolic dataset, which is essential for coherent circuit analysis.
+  • Classification: To address issues of partial inclusion and syntactic ambiguity in constructed syllogism arguments, we employ GPT-4 (OpenAI, 2024) to classify whether arguments contain only truthful premises and whether the middle-terms are syntactically and semantically equivalent.
 This step helps us classify a belief-consistent dataset and a belief-inconsistent dataset.
@@
 B.3 Data Statistics
-We organise data statistics for generated datasets (Table 4).
+For all experiments, we use 90 samples each for both symbolic and non-symbolic arguments to balance analytical depth with computational efficiency.
+This sample size sufficiently yields statistically significant activation patching results ($p<0.05$).
+We organize data statistics for generated datasets (Table 5).
 Statistic SYM BC BI
@@
 Appendix C Embedding Space Analysis Details
-One established method for analysing hidden activation in transformer language models is the logit lens, which projects activation into embedding space (Nostalgebraist, 2020; Elhage et al., 2021; Geva et al., 2022).
-We focus on an input-agnostic approach (Dar et al., 2023; Hanna et al., 2024), utilising both unembedding ($W_{U}$) and embedding ($W_{E}$) matrices to construct a $\mathbb{R}^{|V|\times|V|}$ matrix, where $|V|$ denotes the model’s vocabulary size.
+One established method for analyzing hidden activation in transformer language models is the logit lens, which projects activation into embedding space (Nostalgebraist, 2020; Elhage et al., 2021; Geva et al., 2022).
+We focus on an input-agnostic approach (Dar et al., 2023; Hanna et al., 2024), utilizing both unembedding ($W_{U}$) and embedding ($W_{E}$) matrices to construct a $\mathbb{R}^{|V|\times|V|}$ matrix, where $|V|$ denotes the model’s vocabulary size.
 This study employs the OV circuit (Elhage et al., 2021) of attention heads, formed by attention value and output weights ($W_{V}W_{O}$), to understand how source tokens generally influence output logits.
-The OV circuit-based logit lens for attention head $\mathbf{h}$ is formalised as $W_{E}W^{\mathbf{h}}_{V}W_{O}^{\mathbf{h}}W_{U}\in\mathbb{R}^{|V|\times|V|}$.
-Following the previous works (Elhage et al., 2021; Dar et al., 2023), this formulation omits layer normalisation.
+The OV circuit-based logit lens for attention head $\mathbf{h}$ is formalized as $W_{E}W^{\mathbf{h}}_{V}W_{O}^{\mathbf{h}}W_{U}\in\mathbb{R}^{|V|\times|V|}$.
+Following the previous works (Elhage et al., 2021; Dar et al., 2023), this formulation omits layer normalization.
 Appendix D Circuit Ablation Method Details
@@
 Layer Normalization.
-Before processing by self-attention and MLP processing, each token’s state is normalized to stabilise learning and reduce training time:
+Before processing by self-attention and MLP processing, each token’s state is normalized to stabilize learning and reduce training time:
 $\text{LN}(r_{i}^{l})=\gamma\odot\frac{r_{i}^{l}-\mu_{i}^{l}}{\sqrt{(\sigma_{i}^{l})^{2}+\epsilon}}+\beta$
@@
 Figure 7: Path patching results for (Top) symbolic inputs and non-symbolic inputs (Bottom) in the middle-term corruption setup.
-Note that negative heads are positively influencing heads, as we replace corrupted activation with clean ones in our noising approach.
+Note that negative heads are positively influencing heads, as we replace corrupted activation with clean ones in our noising method.
 Appendix H Localizing Mover Heads
@@
 Red dots indicate attention heads above the threshold ($\tau$) of the patching score.
-Appendix I Non-symbolic Circuit Analysis
-We present the results of our non-symbolic circuit analysis (Figure 9).
-Figure 9: Comprehensive results of the non-symbolic circuit analysis.
-(a) Attention output patching results and (b) residual stream patching results in the middle-term intervention setup.
-(c) OV circuit logit lens results for $\mathbf{h}_{11.10}$, with input and output comprising 26 uppercase letters.
-(d) Residual stream patching results in the all-term corruption setup.
-(e)-(h) Average attention patching weights across the batch: (e) weights for token position $[p]$ attending, and (f)-(h) weights for the last token position attending.
-For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.
-Appendix J Generalizability to Other Syllogisms
+Appendix I Generalizability to Other Syllogisms
 To assess the generalizability of our model, we applied the circuit ablation method to all 15 unconditionally valid syllogisms (Table 4), and the result is shown in Figure 9.
 Unconditionally valid syllogisms are logical arguments that maintain their validity irrespective of the truth values of their premises.
-The validity of these syllogisms is determined solely by their logical form, which is characterised by mood and figure combinations.
+The validity of these syllogisms is determined solely by their logical form, which is characterized by mood and figure combinations.
 The mood of a syllogism is defined by the arrangement of four proposition types ("All (A)", "No (E)", "Some (I)", "Some … not (O)") across its two premises and conclusion.
@@
 Figure 9: Necessity and sufficiency performance results from circuit ablation method for all unconditionally valid syllogistic forms. Labels (e.g., AII-3) denote mood and figure combinations.
-Model S BC BI
-Group 1: Erratic Pattern
-Pythia-70M 0.19 0.02 0.02
-Pythia-160M 0.00 0.12 0.21
-Group 2: Compatible Pattern
-Pythia-410M 0.99 0.40 0.51
-Pythia-1B 0.30 0.31 0.43
-Llama-3.2-1B 1.00 0.81 0.70
-Group 3: Variant Pattern
-Qwen2.5-0.5B 1.00 0.96 0.90
-Qwen2.5-1.5B 1.00 0.99 0.90
-Table 5: Accuracy comparison of different models across three evaluation categories: Symbolic (S), Belief-Consistent (BC), and Belief-Inconsistent (BI).
-Group 1 models exhibit erratic activation patterns, failing to establish a stable reasoning mechanism.
-Group 2 models demonstrate a compatible reasoning circuit with consistent information flow and suppression.
-Group 3, particularly the Qwen models, exhibit a functional but distinct suppression mechanism, where suppression is applied at the final token rather than the $[p]$ position.
-Appendix K Generalizability Across Model Sizes
+Appendix J Generalizability Across Model Sizes
 Our approach is guided by the universality hypothesis, which suggests that similar representational patterns and circuits emerge across models trained on the same dataset (Olah et al., 2020; Kornblith et al., 2019; Li et al., 2015).
@@
 (d) Residual stream patching results in the all-term corruption setup.
-For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.
-Appendix L Generalizability Across Different Models
-Figures 12 and 13 represent the results of applying the same circuit analysis in different model families.
-The $m$-suppression heads for each model were manually selected based on their likelihood, determined by patching scores and spatial positioning within the activation pattern.
-The corresponding model accuracies in the three datasets (Symbolic, Belief-Consistent, and Belief-Inconsistent) are presented in Table 5.
+For clarity, a dash (–) indicates the averaged values for tokens appearing between terms.
+Appendix K Generalizability Across Different Models
+Figures 11 and 12 represent the results of applying the same circuit analysis in different model families.
+The $m$-suppression heads for each model were manually selected based on their likelihood, determined by patching scores and positioning within the activation pattern.
 Figure 11: Comprehensive results of the symbolic circuit analysis across different series of models (Pythia).
@@
 (d) Residual stream patching results in the all-term corruption setup.
-For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.
+For clarity, a dash (–) indicates the averaged values for tokens appearing between terms.
 Figure 12: Comprehensive results of the symbolic circuit analysis across different series of models (Qwen2.5, Llama3.2 and fine-tuned GPT-2 medium).
@@
 (d) Residual stream patching results in the all-term corruption setup.
-For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.
-Appendix M Reproducibility Statement
-Our study is conducted using TransformerLens (Nanda and Bloom, 2022) on an Nvidia A100 GPU with 80GB of memory.
-The dataset and code to reproduce our experiments are available online to encourage future work in the field [^1] http://anonynous_url.com.
-Appendix N Statistical Significance
-For all experiments, we use 90 samples each for both symbolic and non-symbolic arguments to balance analytical depth with computational efficiency.
-This sample size sufficiently yields statistically significant activation patching results ($p<0.05$).
+For clarity, a dash (–) indicates the averaged values for tokens appearing between terms.